### v1

# Linearizations and optimization problems in diffeological spaces

Abstract.
By generalizing the notion of linearization, a concept originally arising from microlocal analysis and symbolic calculus, to diffeological spaces, we make a first proposal setting for optimization problems in this category.
We show how linearizations allow the construction of smooth paths and variational flows without requiring canonical charts or gradients. With these constructions, we introduce a general optimization algorithm adapted to diffeological spaces under weakened assumptions. The method applies to spaces of mappings with low regularity.
Our results show that weak convergence toward minima or critical values can still be achieved under diffeological conditions. The approach extends classical variational methods into a flexible, non-linear infinite-dimensional framework. Preliminary steps to the search for fixed points of diffeological mappings are discussed.
MSC(2020): 49J27, 53C15, 90C53
Keywords: Diffeology, linearization, minimization method
Introduction
The theory of diffeological spaces provides a natural framework for analyzing smooth structures beyond the realm of finite-dimensional manifolds, particularly in contexts where classical differential geometry is insufficient. This flexibility suggests diffeologies to be an essential tool in the study of mapping spaces, infinite-dimensional geometry, and topological vector spaces with weak topologies.
One of the central challenges in applying differential techniques in such settings lies in the absence of canonical local charts or metrics, which are typically used in optimization and variational analysis on manifolds. In this article, we explore and generalize the notion of linearization, originally introduced by Bokobza-Haggiag in the study of pseudo-differential operators on manifolds, and extended later by Widom through symbolic calculus in microlocal analysis. We reinterpret these constructions along the lines of diffeological geometry, and show how they can be leveraged to define optimization algorithms on diffeological spaces. In particular, we introduce a class of minimization methods in the spirit of well-known steepest descent methods, that can be intuitively understood as generalizations of gradient-type methods that do not require explicit knowledge of gradients, but only of a functional and its evaluation along smooth paths.
We then apply this framework to the setting of mapping spaces with low regularity, such as Sobolev spaces $W^{s,p}(M,N)$, where $M$ and $N$ are compact manifolds and $s\in\mathbb{R}$, $1<p<\infty$. We analyze how linearizations of the target manifold $N$ naturally induce linearizations on the corresponding mapping spaces, and establish convergence results for the associated variational algorithm under weak regularity assumptions. This allows us to treat optimization problems in topologies that lack Banach manifold structures, extending classical results into the diffeological domain. A discussion on the construction of invariant spaces, or fixed points, of smooth functions ends the study.
1. Preliminaries on diffeologies
1.1. Diffeologies and Frölicher spaces
In this section, we follow the presentations from [29, 31], and complete the classical exposition [23] along the lines of the recent survey [21].
Definition 1.
Let $X$ be a set.
  • A $p$-parametrization of dimension $p$ on $X$ is a map from an open subset $O\subset\mathbb{R}^{p}$ to $X$.
  • A diffeology on $X$ is a set $\mathcal{P}$ of parametrizations satisfying:
    1. (1)
      All constant maps $\mathbb{R}^{p}\to X$ belong to $\mathcal{P}$.
    2. (2)
      If $\{f_{i}:O_{i}\to X\}_{i\in I}$ is a family of plots in $\mathcal{P}$ compatible on overlaps, and if they define a map $f:\bigcup_{i}O_{i}\to X$, then $f\in\mathcal{P}$.
    3. (3)
      If $f\in\mathcal{P}$ and $g\in C^{\infty}(O^{\prime},O)$, then $f\circ g\in\mathcal{P}$.
A pair $(X,\mathcal{P})$ is called a diffeological space. Given two diffeological spaces $(X,\mathcal{P})$ and $(X^{\prime},\mathcal{P}^{\prime})$, a map $f:X\to X^{\prime}$ is said to be smooth if $f\circ p\in\mathcal{P}^{\prime}$ for every $p\in\mathcal{P}$.
The concept of diffeological space was introduced by J.-M. Souriau [36], inspired by early observations of Chen [6]. A foundational treatment can be found in [23].
Definition 2.
Let $(X,\mathcal{P})$ be a diffeological space. A diffeology $\mathcal{P}^{\prime}$ on $X$ is called a subdiffeology of $\mathcal{P}$ if $\mathcal{P}^{\prime}\subset\mathcal{P}$, or equivalently, if the identity map $\mathrm{Id}_{X}:(X,\mathcal{P}^{\prime})\to(X,\mathcal{P})$ is smooth.
Example 3.
Every nonempty set $X$ carries a minimal diffeology, consisting only of constant maps. This is called the discrete diffeology.
The category of diffeological spaces is extremely broad and includes various pathological examples. To control this, one can work within the more structured category of Frölicher spaces.
Definition 4.
A Frölicher space is a triple $(X,\mathcal{F},\mathcal{C})$ where:
  • $\mathcal{C}$ is a set of curves $c:\mathbb{R}\to X$, called contours.
  • $\mathcal{F}$ is a set of functions $f:X\to\mathbb{R}$ such that $f\circ c\in C^{\infty}(\mathbb{R},\mathbb{R})$ for all $c\in\mathcal{C}$.
  • Conversely, a curve $c:\mathbb{R}\to X$ belongs to $\mathcal{C}$ if and only if $f\circ c\in C^{\infty}(\mathbb{R},\mathbb{R})$ for all $f\in\mathcal{F}$.
A map $f:(X,\mathcal{F},\mathcal{C})\to(X^{\prime},\mathcal{F}^{\prime},\mathcal{C}^{\prime})$ is smooth if and only if $f\circ c\in\mathcal{C}^{\prime}$ and $f^{\prime}\circ f\in\mathcal{F}^{\prime}$ for all $f^{\prime}\in\mathcal{F}^{\prime}$, $c\in\mathcal{C}$.
This notion, introduced in [13], was later formalized in [24]. For an early comparison between diffeological and Frölicher settings, see [28]; see also [1, 29, 30, 31, 37]. In particular, [31] explains that:
Diffeological, Frölicher, and Gâteaux smoothness coincide on Fréchet spaces.
To construct a Frölicher structure on a set $X$, one can proceed as follows: given a set of generating functions $\mathcal{F}_{g}\subset\mathrm{Map}(X,\mathbb{R})$, define
$\mathcal{C}=\{c:\mathbb{R}\to X\mid f\circ c\in C^{\infty}(\mathbb{R},\mathbb{R})\ \forall f\in\mathcal{F}_{g}\},\quad\mathcal{F}=\{f:X\to\mathbb{R}\mid f\circ c\in C^{\infty}(\mathbb{R},\mathbb{R})\ \forall c\in\mathcal{C}\}.$
We say that $\mathcal{F}_{g}$ generates the Frölicher structure $(X,\mathcal{F},\mathcal{C})$.
This structure induces a natural topology on $X$, namely the pullback of the standard topology on $\mathbb{R}$ via the maps in $\mathcal{F}$.
Alternatively, one can begin with a set of generating contours $\mathcal{C}_{g}$, and define:
=$\mathcal{F}=\{f:X\to\mathbb{R}\mid f\circ c\in C^{\infty}(\mathbb{R},\mathbb{R})\ \forall c\in\mathcal{C}_{g}\},\quad\mathcal{C}=\{c:\mathbb{R}\to X\mid f\circ c\in C^{\infty}(\mathbb{R},\mathbb{R})\ \forall f\in\mathcal{F}\}.$
If $X$ is a finite-dimensional smooth manifold, taking $\mathcal{F}=C^{\infty}(X,\mathbb{R})$ and $\mathcal{C}=C^{\infty}(\mathbb{R},X)$ yields a Frölicher structure whose topology matches the standard manifold topology [24]. Moreover, one can define a diffeology on a Frölicher space by:
(1) $\mathcal{P}_{\infty}(\mathcal{F}):=\coprod_{p\in\mathbb{N}^{*}}\left\{f:D(f)\subset\mathbb{R}^{p}\to X\ \middle|\ f\text{ is smooth and }\mathcal{F}\circ f\subset C^{\infty}(D(f),\mathbb{R})\right\}.$
This is known as the nebulae diffeology (cf. [23]).
Proposition 5 ([28]).
Let $(X,\mathcal{F},\mathcal{C})$ and $(X^{\prime},\mathcal{F}^{\prime},\mathcal{C}^{\prime})$ be Frölicher spaces. A map $f:X\to X^{\prime}$ is smooth in the Frölicher sense if and only if it is smooth for the associated diffeologies $\mathcal{P}_{\infty}(\mathcal{F})$ and $\mathcal{P}_{\infty}(\mathcal{F}^{\prime})$.
Hence, we have the chain of implications:
[…]
These implications can be refined and are further analyzed in [37, 28].
Remark 6.
The set of contours $\mathcal{C}$ of a Frölicher space $(X,\mathcal{F},\mathcal{C})$ does not define a diffeology by itself, because a diffeology requires closure under domain restriction. However, one can construct a minimal diffeology $\mathcal{P}_{1}(\mathcal{F})$ consisting of plots locally of the form $c\circ g$, with $c\in\mathcal{C}$ and $g$ a smooth map. Then, Proposition 5 remains valid if we replace $\mathcal{P}_{\infty}$ by $\mathcal{P}_{1}$. This is based on Boman’s theorem [24, p. 26], and detailed in [29, 28, 37].
Definition 7 ([37]).
A diffeology $\mathcal{P}_{\infty}$ defined by a set of functions $\mathcal{F}$ along the lines of (1) is called reflexive. Such diffeologies are those generated by the smooth functions of a Frölicher structure.
We now present additional examples of diffeologies.
Example 8.
Let $k\in\mathbb{N}$ and let $(X,\mathcal{F},\mathcal{C})$ be a Frölicher space. The $C^{k}$-diffeology $\mathcal{P}_{(k)}$ is defined by:
$\forall d\in\mathbb{N}^{*},\ \forall O\subset\mathbb{R}^{d}\text{ open,}\quad\mathcal{P}_{(k)}(O)=\left\{p:O\to X\mid\mathcal{F}\circ p\subset C^{k}(O,\mathbb{R})\right\},$
and we set
$\mathcal{P}_{(k)}=\bigcup_{d\in\mathbb{N}^{*}}\bigcup_{\text{open }O\subset\mathbb{R}^{d}}\mathcal{P}_{(k)}(O).$
This construction generalizes the finite-dimensional case treated in [23].
Proposition 9 ([23, 36]).
Let $(X,\mathcal{P})$ and $(X^{\prime},\mathcal{P}^{\prime})$ be diffeological spaces. The product diffeology on $X\times X^{\prime}$ is defined as the coarsest diffeology for which the projections are smooth. It consists of maps $g:O\to X\times X^{\prime}$ that decompose as $g=(f,f^{\prime})$ with $f\in\mathcal{P}$, $f^{\prime}\in\mathcal{P}^{\prime}$. This definition extends to arbitrary (even uncountable) products.
Proposition 10.
Let $(X,\mathcal{F},\mathcal{C})$ and $(X^{\prime},\mathcal{F}^{\prime},\mathcal{C}^{\prime})$ be Frölicher spaces, with associated diffeologies $\mathcal{P}$ and $\mathcal{P}^{\prime}$. Then $X\times X^{\prime}$ admits a natural Frölicher structure whose set of contours is $\mathcal{C}\times\mathcal{C}^{\prime}$, and whose associated diffeology is the product $\mathcal{P}\times\mathcal{P}^{\prime}$.
This construction also extends to infinite products.
Let $(X,\mathcal{P})$ be a diffeological space and $X^{\prime}$ a set. Given a map $f:X\to X^{\prime}$, the pushforward diffeology $f_{*}(\mathcal{P})$ is the coarsest diffeology on $X^{\prime}$ making $f$ smooth. Conversely, if $(X^{\prime},\mathcal{P}^{\prime})$ is diffeological and $f:X\to X^{\prime}$, the pullback diffeology $f^{*}(\mathcal{P}^{\prime})$ is the finest diffeology on $X$ for which $f$ is smooth.
Proposition 11 ([23, 36]).
Let $(X,\mathcal{P})$ be a diffeological space and $\sim$ an equivalence relation. Then the quotient $X/\sim$ inherits a diffeology $\mathcal{P}/\sim$ defined as the pushforward of $\mathcal{P}$ via the projection $X\to X/\sim$. Locally, plots are of the form $\pi\circ p$, with $p\in\mathcal{P}$.
Let $X_{0}\subset X$ be a subset.
  • If $X$ is diffeological with $\mathcal{P}$, the subset diffeology on $X_{0}$ is the restriction $\mathcal{P}_{0}=\{p\in\mathcal{P}\mid\mathrm{Im}(p)\subset X_{0}\}$.
  • If $(X,\mathcal{F},\mathcal{C})$ is Frölicher, we define the induced structure on $X_{0}$ using the restrictions of maps in $\mathcal{F}$ and contours in $\mathcal{C}$ with image in $X_{0}$.
Example 12.
Let $X,Y$ be sets and consider the diffeological space $Y^{X}$ with the functional product diffeology, defined as the largest diffeology making all evaluation maps $\mathrm{ev}_{x}:Y^{X}\to Y$ smooth. Any subspace of $Y^{X}$ admitting this property inherits a subdiffeology of this functional product diffeology.
Proposition 13 ([23]).
Let $X,Y,Z$ be diffeological spaces. Then the canonical identifications
$C^{\infty}(X\times Y,Z)\simeq C^{\infty}(X,C^{\infty}(Y,Z))\simeq C^{\infty}(Y,C^{\infty}(X,Z))$
hold, where all function spaces are equipped with the functional diffeology.
1.2. Groups, vector spaces, algebra and fiber bundles in diffeologies
Given an algebraic structure, one can naturally define a compatible diffeological or Frölicher structure. For instance, see [26, 29], and more specifically [23, pp. 66–68] for the case of vector spaces.
If $\mathbb{R}$ is equipped with its standard diffeology (or Frölicher structure), then an $\mathbb{R}$-vector space $V$ equipped with a diffeology (or Frölicher structure) is called a diffeological (resp. Frölicher) vector space if vector addition and scalar multiplication are smooth maps.
Definition 14.
Let $G$ be a group endowed with a diffeology (resp. a Frölicher structure). Then $G$ is called a diffeological (resp. Frölicher) group if both the multiplication map $G\times G\to G$ and the inversion map $G\to G$ are smooth.
This notion applies similarly to rings, modules, algebras, and other algebraic structures where compatibility with the diffeological or Frölicher framework requires the corresponding algebraic operations to be smooth.
Moreover, one can define fiber bundles in the diffeological context. A diffeological fiber bundle is a surjective smooth map $\pi:E\to B$ between diffeological spaces such that locally (in the diffeological sense), $E$ is diffeomorphic to $B\times F$, with $F$ the typical fiber, and the transition functions are smooth in the functional diffeology sense.
Many classical constructions—direct sums, duals, subspaces—have analogs in the diffeological category, but must be handled carefully since the lack of local triviality (in the usual topological sense) may affect their behavior. In particular, when dealing wit tangent and cotangent spaces often belong to extended classes called pseudo-bundles (see e.g. [34]), namely,
  • the cotangent space of a diffeological space is defined in a widely accepted way in the literature; but it appears that there is no typical fiber. The fivbers are vector spaces, but they can be no-isomorphic. For example, the cross has a cotangent bundle with fibers of dimension 1 or 2. This is called a vector pseudo bundle
  • the notion of tangent bundle is even more difficult to study because many non-equivalent versions are actually proposed, and each of them has its own motivations.
For an overview on tangent spaces, cotangent spaces, vector pseudo bundles, we refer to [21].
2. Linearizations from Finite-Dimensional Manifolds to Diffeological Spaces and Infinite-Dimensional Settings
2.1. Linearizations of a Manifold
Bokobza-Haggiag introduces the notion of linearization in [4], aiming to extend local analytical constructions—such as differential operators—into a global structure on a smooth manifold $M$. Let us first recall the classical concept:
Definition 15.
A local linearization at a point $x\in M$ is a diffeomorphism
$\psi_{x}:U\subset M\longrightarrow V\subset\mathbb{R}^{n},$
such that $\psi_{x}(x)=0$ and $d\psi_{x}|_{x}:T_{x}M\to\mathbb{R}^{n}$ is the identity (under a fixed identification).
The paper [4] generalizes this to a global notion:
Definition 16 ([4]).
A (global) linearization is a smooth map $\nu:X\times X\rightarrow TX$ such that:
  1. (1)
    Let $p_{1}:X\times X\rightarrow X$ denote the projection onto the first component, and let $\pi:TX\rightarrow X$ denote the tangent bundle projection. Then the following diagram commutes:
    ${X\times X}$${TX}$${X}$$\scriptstyle{\nu}$$\scriptstyle{p_{1}}$$\scriptstyle{\pi}$
  2. (2)
    For all $x\in X$, one has $\nu(x,x)=0$.
  3. (3)
    For all $x\in X$, the differential $D\nu(x,\cdot)|_{y=x}:T_{x}X\to T_{0}(T_{x}X)\simeq T_{x}X$ is the identity.
Using classical geometric techniques, one can derive the following properties:
  1. (1)
    By the inverse function theorem, a linearization defines a local linearization at each point $x\in X$.
  2. (2)
    There exists a neighborhood $\Omega$ of the diagonal $\Delta(X)\subset X\times X$ such that $\nu|_{\Omega}$ is a diffeomorphism onto a neighborhood of the zero section of $TX$.
  3. (3)
    The map
    $l_{x}^{y}:\xi\in T^{*}X\mapsto d_{y}\langle\nu(x,y),\xi\rangle$
    is a smooth linear map from $T^{*}_{y}X$ to $T^{*}_{x}X$.
  4. (4)
    Every manifold $X$ admits a linearization.
  5. (5)
    A manifold $X$ admits an atlas with affine transition functions if and only if there exists a linearization $\mu$ of $X$ such that, for all $(x,y,z)\in X^{3}$ with $(x,y),(y,z),(x,z)\in\Omega$, one has
    $l_{x}^{z}=l_{y}^{z}\circ l_{x}^{y}.$
We present an equivalent formulation of Definition 16 that will be helpful for further generalization.
Proposition 17.
The following condition is equivalent to Definition 16.
Define $k:T^{*}X\times X\to\mathbb{R}$ by
$k(\xi,y):=\langle\nu(x,y),\xi\rangle,\quad\text{with }x=\pi(\xi).$
Then $\nu$ is a global linearization if and only if:
  1. (1)
    $k(\xi,x)=0$ for all $\xi\in T^{*}X$ (where $x=\pi(\xi)$);
  2. (2)
    The differential $d_{y}k(\xi,y)|_{y=x}:T^{*}_{x}X\to T^{*}_{x}X$ is the identity.
2.2. Exponential Map and Linearization on a Compact Riemannian Manifold
Let $(M,g)$ be a compact, connected, orientable, boundaryless Riemannian manifold. By the Hopf–Rinow theorem, $M$ is geodesically complete, so the exponential map
$\exp_{p}:T_{p}M\to M$
is defined on the entire tangent space $T_{p}M$ for every $p\in M$.
2.2.1. Properties of the exponential map
We follow classical references such as [3, 14]. Let $v\in T_{p}M$ and consider the geodesic $\gamma(t)=\exp_{p}(tv)$. Then $d\exp_{p}|_{v}$ satisfies:
  • Identity at the origin:
    $d\exp_{p}|_{0}=\mathrm{id}_{T_{p}M}.$
  • Critical and regular points: A vector $v\in T_{p}M$ is a regular point if $d\exp_{p}|_{v}$ is invertible. Otherwise, $v$ is critical, and $\exp_{p}(v)$ is a conjugate point along the geodesic.
  • Local diffeomorphism: There exists $\varepsilon>0$ such that $\exp_{p}$ is a diffeomorphism from $B(0,\varepsilon)\subset T_{p}M$ onto its image.
  • Injectivity radius: The injectivity radius at $p$ is the supremum of $r>0$ such that $\exp_{p}$ is a diffeomorphism on $B(0,r)$. For compact $M$, this radius is uniform and strictly positive.
  • Uniform boundedness: If the Riemann curvature is uniformly bounded, then the norm of $d\exp_{p}$ can be estimated via comparison theorems (e.g., Rauch’s). There exists a uniform radius $r>0$ such that $\exp_{p}$ is a diffeomorphism on $B(0,r)\subset T_{p}M$.
2.2.2. The Exponential Map Defines a Linearization in the Sense of Bokobza-Haggiag
Let $M$ be as above. Since the injectivity radius $r>0$ is uniform, the exponential map defines a diffeomorphism:
$\exp:B_{r}=\left\{v\in TM\ \middle|\ \langle v,v\rangle<r^{2}\right\}\to V_{r}\subset M\times M,$
where $V_{r}$ is a neighborhood of the diagonal $\Delta(M)$ in $M\times M$.
Define $\delta:TM\to V_{r}$ by
$\delta(v):=\left(x,\exp_{x}\left(\frac{r}{\sqrt{1+\langle v,v\rangle}}v\right)\right),$
where $v\in T_{x}M$.
Then:
  • $\delta\in C^{\infty}(TM,M\times M)$;
  • For any linearization $\nu$ of $M$ in the sense of Bokobza-Haggiag, we have:
    $D_{0}\left(\left(\frac{1}{r}\nu\right)\circ\delta\right)=\mathrm{Id}_{T_{x}X}=D_{x}\left(\delta\circ\left(\frac{1}{r}\nu\right)\right),$
    where $D_{0}$ is the dervative along the fibers of $T_{X}$ and $D_{x}$ is the derivative along the second variable of $X\times X.$
Using the Riemannian metric, we identify $TM\cong T^{*}M$. Hence, both tangent and cotangent bundles may be used interchangeably in the characterization of linearizations.
These properties will serve as a reference point when generalizing linearizations to diffeological spaces.
3. Extending Classical Optimization Methods in Linearizations
3.1. A Generalized Notion of Linearization
Definition 18.
Let $(X,\mathcal{P})$ be a diffeological space. Let $E$ be a diffeological vector pseudo-bundle over $X$ with projection $\pi_{E}$, not necessarily equipped with local trivializations or smooth partitions of unity, but such that the inclusion map $T^{*}X\subset E$ is amorphism of diffeological pseudo-bundles. A linearization of $X$ is a pair $(\nu,\delta)$ of smooth maps satisfying:
  1. (1)
    $\delta:E\rightarrow X\times X$ is smooth,
  2. (2)
    $\delta^{-1}_{|T^{*}X}(\Delta(X))$ is the zero section of $T^{*}X$, and the diagram
    ${E}$${X\times X}$${X}$$\scriptstyle{\delta}$$\scriptstyle{\pi_{E}}$$\scriptstyle{p_{1}}$
    commutes,
  3. (3)
    $\nu:X\times X\rightarrow E$ is smooth,
  4. (4)
    $\forall(x,y)\in X^{2},$ we have $\nu(x,y)=0\Leftrightarrow x=y$,
  5. (5)
    $\forall x\in X,$ the derivative $d_{x}\nu(x,.)\neq 0$,
  6. (6)
    The diagram
    ${X\times X}$${E}$${X}$$\scriptstyle{\nu}$$\scriptstyle{p_{1}}$$\scriptstyle{\pi_{E}}$
    commutes,
  7. (7)
    For every $x\in X$, the induced map from $T^{*}X$ to $T^{*}X$ by $(\delta\circ\nu)(x,.)\in C^{\infty}(X,X)$ restricts to an injective map on $T^{*}_{x}X$.
This definition mirrors the notation in Section 2.2.2 to highlight differences with the classical definition by Bokobza-Haggiag. Notably, condition 4 does not generally hold in the classical case when $\nu$ targets $TX$ or $T^{*}X$. The non-trivial homotopy type of $X$ is a main obstruction such a global property.
This technical obstacle is circumvented by enlarging the target space to a pseudo-bundle $E$ over $X$ containing $T^{*}X$ as a sub-pseudo-bundle. The main motivation is that $T^{*}X$ is well-defined in the diffeological setting, unlike the tangent bundle, for which several constructions exist. We now verify the existence of such linearizations in standard cases:
Theorem 19.
Let $M$ be a compact smooth manifold. Then there exists a linearization in the sense of Definition 18 with $E=T^{*}M\times\mathbb{R}$.
Proof.
[…]
Define the map
$(x,y)\mapsto\left(g\circ\nu_{BH}(x,y),d(x,y)\right)\in T^{*}M\times\mathbb{R}.$
This map is continuous and can be approximated by a smooth one $\nu$ via Whitney’s density theorem. Then, define:
$\delta(v,k)=\exp\left(\frac{r}{\sqrt{1+\langle(1+k)v,(1+k)v\rangle}}v\right),$
where $r$ is the uniform injectivity radius of $M$. This yields a linearization in the required sense. ∎
3.2. Distance Induced by a Linearization
We now introduce an intrinsic structure for measuring convergence on $X$ using the linearization $(\delta,\nu)$. Note that the map $\delta$ plays a central role in this construction.
Assumption D.
The pseudo-bundle $E$ is equipped with a function $gap:E\rightarrow\mathbb{R}$ that satisfies:
  • The restriction $gap\in C^{\infty}(E\setminus X,\mathbb{R})$, where $X$ is identified with the zero section;
  • $gap(v)=0\iff v=0$ in the fiber $E_{x}$ over $x=\pi_{E}(v)$;
  • $gap$ is continuous for the $D$-topology and bounded on $X$.
This function defines a basis of neighborhoods of $X$ (seen as the zero section). It can be understood as a diffeological adaptation of teh notion of uniform structure111we recall that, in general topology, a uniform structure on a topological space $X$ is a basis of neighborhoods of the diagonal $\Delta(X)$ of $X\times X.$ Within additional technical assumptions, it enables to develop a notion of Cauchy sequences without prescribed metric. on $X$ and allows us to define:
Definition 20.
Let $(x,y)\in X\times X$ and $(u,v)\in E\times E$. We define:
  • $d_{X}(x,y)=\sup_{z\in X}\left|gap(\nu(z,x))-gap(\nu(z,y))\right|$;
  • If $\pi_{E}(u)=\pi_{E}(v)$, then
    $d_{E}(u,v)=d_{X}(\delta(u-v))+d_{X}(\delta(v-u)).$
To simplify notation, subscripts are dropped when context permits.
Theorem 21.
Let $X$ be a diffeological space equipped with a linearization $(\delta,\nu)$ and a gap function satisfying Assumption D. Then $d_{X}$ is a distance on $X$ and $d_{E}$ is a distance on each fiber of $E$.
Proof.
We verify the axioms of a metric for $d_{X}$ and $d_{E}$:
  1. (1)
    Separating property:
    $gap(\nu(x,y))\neq 0\iff x\neq y.$
    Thus, $d_{X}(x,x)=0$, and if $x\neq y$, then choosing $z=x$ gives $d_{X}(x,y)>0$.
    Similarly, $d_{E}(u,v)==0$ if and only if $u=v$.
  2. (2)
    Symmetry: immediate from definitions.
  3. (3)
    Triangle inequality: For $x,y,z\in X$ and all $w\in X$:
    $\left|gap(\nu(w,x))-gap(\nu(w,z))\right|$$\leq$ $\left|gap(\nu(w,x))-gap(\nu(w,y))\right|$
    $+\left|gap(\nu(w,y))-gap(\nu(w,z))\right|.$
    Taking suprema yields $d_{X}(x,z)\leq d_{X}(x,y)+d_{X}(y,z)$.
Remark 22.
Convexity of the gap function is not required for these properties.
Definition 23.
Let $\tilde{X}$ denote the metric completion of $X$ with respect to $d_{X}$.
4. Gradient and steepest descent methods… (Almost) Without gradient
Let us now consider a minimization problem of a given functional $F:X\rightarrow\mathbb{R}$ or an equation $F(u)=0$ on a diffeological space [^2] We recall that solving $F(x)=0$ is equivalent to solving the minimization problem for $|F|.$. We define the following basic notion:
Definition 24.
A method $\mathcal{M}$ is a mapping that depends on:
  • a smooth function $f\in C^{\infty}(I;\mathbb{R})$, where $I\subset\mathbb{R}$ is compact,
  • and a point $x\in I$,
and produces a point $\mathcal{M}_{f}(x)\in I$ such that:
  1. (1)
    $\forall x\in I,\quad f(\mathcal{M}_{f}(x))\leq f(x)$,
  2. (2)
    $\forall x\in I,\quad f(\mathcal{M}_{f}(x))=f(x)\iff\mathcal{M}_{f}(x)=x$,
  3. (3)
    $\mathcal{M}_{f}$ is smooth on any subinterval $J\subset I$ where $df\neq 0$.
By this definition, we intend to define a method of the type of the gradient methods, Newton methods for convex functions, and more generally any iterative method that (often approximatively at a finite number of iterations) intend to produce the minimum of a smooth function $f$ of prescribed type on a compact interval.
4.1. The Algorithm Associated to a Method $\mathcal{M}$
To provide context, we first sketch the general principle of the method. The core idea is to build a smooth curve $\gamma$ on $X$, replacing the gradient direction used in classical optimization by a path derived from the linearization.
This path is defined by the linearization map $\delta$, such that:
$\gamma(t)=\delta(t\,d_{x}F).$
Data: Smooth function $F:X\to\mathbb{R}$, initial point $x_{0}$, tolerance $\varepsilon>0$, maximum iterations $N_{\max}$.
Result: Approximate solution to $F(u)=0$.
begin
       $x\leftarrow x_{0}$
       $\gamma(t)\leftarrow\delta(t\,d_{x}F)$
       $d\leftarrow 1+\varepsilon$
       for $n\leftarrow 1$ to $N_{\max}$ do
             if $d<\varepsilon$ then
                   return Approximate zero of the gradient found at $x$
            if $d_{x}F=0$ then
                   return Exact critical point found at $x$
            $d\leftarrow x-\mathcal{M}_{F\circ\gamma}(x)$
             $x\leftarrow\mathcal{M}_{F\circ\gamma}(x)$
             $\gamma(t)\leftarrow\delta(t\,d_{x}F)$
      return Warning: Maximum number of iterations reached.
Algorithm 1 Gradient method with generalized linearization
Remark 25.
Our algorithm is an adaptation of the algorithm presented in [18, 19, 20], where the authors use a retraction map from $TX$ to $X$ in place of $\delta$, and replace the gradient by the differential $df$ in $T^{*}X$.
4.2. Classical A Priori Convergence Results to Weak Solutions
To analyze convergence, we make the following assumption:
  Assumption S.
There exists a sequence $(x_{n})\in X^{\mathbb{N}}$ defined by:
$\left\{\begin{array}[]{l}x_{0}\text{ given in }X,\\ \gamma_{n}(t)=\delta(t\,d_{x_{n}}F),\\ x_{n+1}=\mathcal{M}_{F\circ\gamma_{n}}(x_{n}).\end{array}\right.$
We ignore the iteration bound and set $\varepsilon=0$. If for some $n$, $d_{x_{n}}F=0$, then the sequence becomes stationary at $x_{n}$. We additionally assume that: - $X$ is topologically embedded in a complete metric vector space $Y$, - $X$ has compact closure in $Y$, - and $F$ extends continuously to $\overline{X}\subset Y$.
Theorem 26.
Let $S_{Y}$ be the set of limit points of the sequence $(x_{n})$. Then:
  • $F$ is constant on each connected component of $S_{Y}$,
  • and $-\infty<F(S_{Y})\leq F(x_{0})$.
Proof.
Compactness of $\overline{X}$ implies $S_{Y}\neq\emptyset$. Let $y\in S_{Y}$. Then there exists an injective extraction map $\epsilon:\mathbb{N}\to\mathbb{N}$ such that $x_{\epsilon(n)}\to y$.
From Definition 24, $F(x_{n+1})\leq F(x_{n})$, hence $-\infty<F(y)=\lim F(x_{\epsilon(n)})$.
If $y^{\prime}\in S_{Y}$ is another limit point, there exists another injection $\epsilon^{\prime}:\mathbb{N}\to\mathbb{N}$ with $x_{\epsilon^{\prime}(n)}\to y^{\prime}$. For every $n$, there exists $m(n)$ such that $\epsilon^{\prime}(m(n))>\epsilon(n)$, so
$F(y^{\prime})=\lim F(x_{\epsilon^{\prime}(m(n))})\leq\lim F(x_{\epsilon(n)})=F(y).$
Thus $|F(y)|=|F(y^{\prime})|$. Since $F$ is continuous, $F$ is constant on any connected component of $S_{Y}$. ∎
4.3. Convexity Assumptions and Strong Convergence
We now study the relationship between $S_{Y}$ and minimizers of $F$. The set $S_{Y}$ may be interpreted as weak solutions to $F(u)=0$, while $X\cap S_{Y}$ corresponds to strong solutions, along the lines of the study [33] in diffeological spaces. The space $Y$ provides the necessary structure to study convergence, as diffeological spaces lack intrinsic topologically complete distances.
We first recall the following:
Lemma 27.
Let $P:C\to X$ be a smooth injective map from a convex set $C$ that extends to an injective map $\overline{P}:\overline{C}\to\overline{X}$. Then $\overline{C}$ is convex, and $F\circ\overline{P}$ is convex on $\overline{C}$.
Proof.
Let $x,y\in\overline{X}$, and $\lambda\in[0,1]$. There exist sequences $(c_{n}),(d_{n})\subset C$ with $P(c_{n})\to x$, $P(d_{n})\to y$, and thus:
$P(\lambda c_{n}+(1-\lambda)d_{n})\to\lambda x+(1-\lambda)y.$
Since $C$ is convex, the convex combination lies in $C$, and the limit is in $\overline{X}$, proving that $\overline{C}$ is convex.
Now, by convexity of $F$ on $C$:
$F(P(\lambda c_{n}+(1-\lambda)d_{n}))\leq\lambda F(P(c_{n}))+(1-\lambda)F(P(d_{n})).$
Taking limits:
$F(\lambda x+(1-\lambda)y)\leq\lambda F(x)+(1-\lambda)F(y).$
Theorem 28.
If $F$ is strictly convex on $Im(\overline{P})$ and $P$ satisfies Lemma 27, then $S_{Y}$ contains a unique point.
5. Application to Mapping Spaces with Low Regularity
The notion of manifolds of smooth maps used here follows the framework in [11]; see also [24], [12]. Let $E$ be a smooth vector bundle over a compact, connected, boundaryless manifold $M$ of dimension $m$. Denote by $C^{\infty}(M,E)$ the space of smooth sections $\sigma$ of $E$.
Let ${\mathbb{K}}=\mathbb{R}$ or $\mathbb{C}$, and let $N$ be a smooth Riemannian manifold without boundary. Then:
- $C^{\infty}(M,N)$ denotes the set of smooth maps $f:M\rightarrow N$,
- $C^{\infty}(M,G)$ denotes the space of smooth maps from $M$ to a Lie group $G$.
The space $C^{\infty}(M,N)$ is a smooth Fréchet manifold, and $C^{\infty}(M,G)$ is a smooth Fréchet Lie group.
The connected components of $C^{\infty}(M,N)$ correspond to the homotopy classes of based maps $f:M\rightarrow N$. Note that two such components need not be modeled on the same topological vector space. However, if $f$ and $g$ lie in the same path-connected component, then the bundles $f^{*}TN$ and $g^{*}TN$ are isomorphic. The atlas structure of $C^{\infty}(M,N)$ can be found e.g. in [27].
5.1. Spaces of Mappings with Low Regularity
[…]
We follow and generalize [32, 21]:
Definition 29.
Let $(s,p)\in\mathbb{R}\times(1,+\infty)$, let $M$ be a compact Riemannian manifold without boundary, and let $N$ be a Riemannian submanifold embedded in $\mathcal{M}$. Then, the space $W^{s,p}(M,N)$ is defined as the completion of $C^{\infty}(M,N)$ in $W^{s,p}(M,\mathcal{A})$.
We emphasize that the classical condition $s<\dim(M)/p$ is not assumed here. When $s\leq\dim(M)/p$, this results in a weaker differential structure on $W^{s,p}(M;\mathcal{A})$, as detailed in [32].
Proposition 30.
The space $W^{s,p}(M,N)$ is a diffeological space equipped with the subset diffeology inherited from $W^{s,p}(M,\mathcal{A})$.
This completion coincides with the $\overline{X}$ construction used in earlier sections.
5.2. From Linearizations on $N$ to Linearizations on $C^{\infty}(M,N)$ and the Equation $F(u)=0$ in $W^{s,p}(M,N)$
Since $N$ is compact, it admits a linearization $(\delta,\nu)$ between $TN\times\mathbb{R}$ and $N\times N$.
Definition 31.
Let $f\in C^{\infty}(M,N)$. Then:
  • For all $g\in C^{\infty}(M,N)$, define $\nu(f,g):x\mapsto\nu(f(x),g(x))\in f^{*}(T^{*}N)\times\mathbb{R}$.
  • For any section $s$ of $f^{*}(T^{*}N)\times\mathbb{R}$, define $\delta(s)=(f,g)$ such that for all $x\in M$, $(f(x),g(x))=\delta(s(x))$.
Proposition 32.
The maps
$\delta\in C^{\infty}(C^{\infty}(M,TN\times\mathbb{R}),C^{\infty}(M,N)\times C^{\infty}(M,N)),\quad\nu\in C^{\infty}(C^{\infty}(M,N)\times C^{\infty}(M,N),C^{\infty}(M,TN\times\mathbb{R}))$
define a linearization of $C^{\infty}(M,N)$.
We now consider a functional equation:
$F(u)=0,\quad\text{with }F:D\subset W^{s,p}(M,N)\rightarrow\mathbb{R},$
where $D\cap C^{\infty}(M,N)$ is dense in $D$, and $D$ is closed in $W^{s,p}(M,N)$. We assume that:
 Assumption R. The linearization $(\delta,\nu)$ on $C^{\infty}(M,N)$ restricts to a linearization on the subset $D\cap C^{\infty}(M,N)$ equipped with the subset diffeology.
This assumption is non-trivial in general. For instance, if $D\cap C^{\infty}(M,N)$ inherits the discrete diffeology, then $T^{*}C^{\infty}(M,N)=C^{\infty}(M,N)$ and the cotangent space has trivial fibers. This highlights the potential limitations of such restrictions, and suggests that constructing a dedicated linearization directly on $D\cap C^{\infty}(M,N)$ may be more effective for algorithmic implementations.
In what follows, we restrict ourselves to the case $D=W^{s,p}(M,N)$ and $D\cap C^{\infty}(M,N)=C^{\infty}(M,N)$.
Theorem 33.
Let $s\leq 0$, and define a sequence $(x_{n})$ inductively by:
$\left\{\begin{array}[]{l}x_{0}\in C^{\infty}(M,N)\text{ is prescribed},\\ \gamma_{n}(t)=\delta(t\,d_{x_{n}}F),\\ t_{n}\in[-\epsilon,\epsilon]\text{ minimizes }F\circ\gamma_{n}(t),\\ x_{n+1}=\gamma_{n}(t_{n}).\end{array}\right.$
Then:
  • The sequence $(x_{n})$ admits a non-empty adherence set in $W^{s,p}(M,N)$.
  • If $(x_{n})$ lies in the image of a smooth parametrization $P:C\rightarrow W^{s,p}(M,N)$ where $C$ is convex and $F\circ P$ is strictly convex on $C$, then $(x_{n})$ converges to $x\in W^{s,p}(M,N)$ such that $|F(x)|$ is minimal on $\operatorname{Im}(P)$.
Proof.
Let $Y=W^{s,p}(M,\mathcal{A})$. Since $s\leq 0$, we have the continuous inclusion $C^{0}(M,\mathcal{A})\subset W^{s,p}(M,\mathcal{A})$. Moreover, since $C^{\infty}(M,N)$ is bounded in $C^{0}(M,N)$, its closure $W^{s,p}(M,N)$ is compact in $W^{s,p}(M,\mathcal{A})$. The result then follows from Theorems 26 and 28. ∎
6. In search for fixed points… if possible
Let us now consider the fixed point problem
(2) $f(x)=x$
for a smooth function $f:X\rightarrow X$ defined on a diffeological space $X$ with linearization $\delta,\nu)$ and gap function $gap$. Then Problem (2) is equivalent to the minimization problem of the function
(3) $F:x\in X\mapsto gap\circ\nu(x,f(x)).$
By the construction of the functions $gap,$ $\nu$ and $f,$ we get easily:
Lemma 34.
Let $fix(f)$ be the set of fixed points of $f$ in $X.$ Then, the function $F$ is smooth on $X-fix(F).$
Let us now make necessarily heuristic considerations, because of the very general setting that we consider.
  1. a.
    If $X$ is compactly embedded in a complete topological space $Y,$ and if $F$ extends by density to a function $\overline{F}$ defined on the completion $\overline{X}$ of $X$ in $Y,$ then, for any starting point $x_{0}\in X,$ the set $S_{Y}$ defined in Theorem 26 is not empty and this is a set where $\overline{F}$ is constant. Moreover, $\forall x\in S_{Y}\cap X,$ $\exists n\in\mathbb{N}^{*},$ $f^{n}(x)\in S_{Y}\cap X.$
  2. b.
    We also can remark that
    $\forall x\in X,\quad 0\leq F(x)\leq d_{X}(x,f(x)).$
    Therefore, if the sequence $(x_{n})$ converges to a fixed point of $f$ for $d_{X},$ this is a sufficient condition to ensure that $limF(x_{n})=0$ even if $F$ is not smooth at fixed points of $f.$
  3. c.
    If moreover $f$ extends to a continuous function $\tilde{f}$ on the completion $\tilde{X}$ of $X$ for $d_{X},$ and if this completion is compact, the adherence set $S_{\tilde{X}}$ is an invariant set for $\tilde{f}.$
Therefore, we can hope that the framework of linearizations can provide new insights for the generalization of classical techniques in fixed point theory and discrete dynamical systems on diffeological spaces. However, the full study of precise examples require techniques that are out of the scope of the present paper that only consider more abstract aspects of these problems.
Outlook
We have presented a abstract framework for variational methods in diffeological spaces based on the concept of linearization, thereby extending classical optimization strategies beyond the confines of smooth Banach or Hilbert manifolds. By abstracting the role of the gradient through a diffeological notion of directional variation, we introduced a flexible algorithm capable of operating in spaces with low or non-standard regularity.
The linearizations from the works of Bokobza-Haggiag and Widom were shown to provide complementary tools for defining smooth paths and comparisons in spaces that may lack coordinate charts or Riemannian metrics. The proposed generalization in the context of diffeological spaces enabled us to design a approximation method for minimization methods.
Several directions remain open, including the intrinsic construction of linearizations for diffeological subspaces, the analysis of convergence rates, and the extension to constraints or flows defined on stratified or singular spaces, and certainly the most important question will be the “concrete” application of the principles described here into non trivial examples. We hope that this work contributes to bridging the gap between differential geometry and optimization in generalized smooth settings and ou future efforts will remain on the applications of these techniques to infinite dimensional problems for functions with low regularity.
Acknowledgements:
J.-P.M was partially supported by ANID (Chile) via the FONDECYT grant #1201894 during the first stages of this work. He also thanks the France 2030 framework programme Centre Henri Lebesgue ANR-11-LABX-0020-01 for creating an attractive mathematical environment.
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work the author used ChatGPT in order to smoothen the expression in English. After using this tool/service, the author reviewed and edited the content as needed and takes full responsibility for the content of the publication.
Data availability statement
No dataset is related to this work.