### v2

# Linearizations and optimization problems in diffeological spaces

Abstract.
By generalizing the notion of linearization, a concept originally arising from microlocal analysis and symbolic calculus, to diffeological spaces, we make a first proposal setting for optimization problems in this category.
We show how linearizations allow the construction of smooth paths and variational flows without requiring canonical charts or gradients. With these constructions, we introduce a general optimization algorithm adapted to diffeological spaces under weakened assumptions. The method applies to spaces of mappings with low regularity.
Our results show that weak convergence toward minima or critical values can still be achieved under diffeological conditions. The approach extends classical variational methods into a flexible, non-linear infinite-dimensional framework. Preliminary steps to the search for fixed points of diffeological mappings are discussed.
MSC(2020): 49J27, 53C15, 90C53
Keywords: Diffeology, linearization, minimization method
Introduction
The theory of diffeological spaces provides a natural framework for analyzing smooth structures beyond the realm of finite-dimensional manifolds, particularly in contexts where classical differential geometry is insufficient. This flexibility suggests diffeologies to be an essential tool in the study of mapping spaces, infinite-dimensional geometry, and topological vector spaces with weak topologies.
One of the central challenges in applying differential techniques in such settings lies in the absence of canonical local charts or metrics, which are typically used in optimization and variational analysis on manifolds. See also [8, 1]. In this article, we explore and generalize the notion of linearization, originally introduced by Bokobza-Haggiag in the study of pseudo-differential operators on manifolds, and extended later by Widom through symbolic calculus in microlocal analysis. [51] [7] We reinterpret these constructions along the lines of diffeological geometry, and show how they can be leveraged to define optimization algorithms on diffeological spaces. In particular, we introduce a class of minimization methods in the spirit of well-known steepest descent methods, that can be intuitively understood as generalizations of gradient-type methods that do not require explicit knowledge of gradients, but only of a functional and its evaluation along smooth paths.
We then apply this framework to the setting of mapping spaces with low regularity, such as Sobolev spaces $W^{s,p}(M,N)$, where $M$ and $N$ are compact manifolds and $s\in\mathbb{R}$, $1<p<\infty$. We analyze how linearizations of the target manifold $N$ naturally induce linearizations on the corresponding mapping spaces, and establish convergence results for the associated variational algorithm under weak regularity assumptions. This allows us to treat optimization problems in topologies that lack Banach manifold structures, extending classical results into the diffeological domain. A discussion on the construction of invariant spaces, or fixed points, of smooth functions ends the study.
1. Preliminaries on diffeologies
1.1. Diffeologies and Frölicher spaces
In this section, we follow the presentations from [37, 39], and complete the classical exposition [31] along the lines of the recent survey [29].
Definition 1.
Let $X$ be a set.
  • A $p$-parametrization of dimension $p$ on $X$ is a map from an open subset $O\subset\mathbb{R}^{p}$ to $X$.
  • A diffeology on $X$ is a set $\mathcal{P}$ of parametrizations satisfying:
    1. (1)
      All constant maps $\mathbb{R}^{p}\to X$ belong to $\mathcal{P}$.
    2. (2)
      If $\{f_{i}:O_{i}\to X\}_{i\in I}$ is a family of plots in $\mathcal{P}$ compatible on overlaps, and if they define a map $f:\bigcup_{i}O_{i}\to X$, then $f\in\mathcal{P}$.
    3. (3)
      If $f\in\mathcal{P}$ and $g\in C^{\infty}(O^{\prime},O)$, then $f\circ g\in\mathcal{P}$.
A pair $(X,\mathcal{P})$ is called a diffeological space. Given two diffeological spaces $(X,\mathcal{P})$ and $(X^{\prime},\mathcal{P}^{\prime})$, a map $f:X\to X^{\prime}$ is said to be smooth if $f\circ{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}c}\in\mathcal{P}^{\prime}$ for every $p\in\mathcal{P}$.
The concept of diffeological space was introduced by J.-M. Souriau [46], inspired by early observations of Chen [11]. A foundational treatment can be found in [31].
Definition 2.
Let $(X,\mathcal{P})$ be a diffeological space. A diffeology $\mathcal{P}^{\prime}$ on $X$ is called a subdiffeology of $\mathcal{P}$ if $\mathcal{P}^{\prime}\subset\mathcal{P}$, or equivalently, if the identity map $\mathrm{Id}_{X}:(X,\mathcal{P}^{\prime})\to(X,\mathcal{P})$ is smooth.
Example 3.
Every nonempty set $X$ carries a minimal diffeology, consisting only of constant maps. This is called the discrete diffeology.
The category of diffeological spaces is extremely broad and includes various pathological examples. To control this, one can work within the more structured category of Frölicher spaces.
Definition 4.
A Frölicher space is a triple $(X,\mathcal{F},\mathcal{C})$ where:
  • $\mathcal{C}$ is a set of curves $c:\mathbb{R}\to X$, called contours.
  • $\mathcal{F}$ is a set of functions $f:X\to\mathbb{R}$ such that $f\circ c\in C^{\infty}(\mathbb{R},\mathbb{R})$ for all $c\in\mathcal{C}$.
  • Conversely, a curve $c:\mathbb{R}\to X$ belongs to $\mathcal{C}$ if and only if $f\circ c\in C^{\infty}(\mathbb{R},\mathbb{R})$ for all $f\in\mathcal{F}$.
A map $f:(X,\mathcal{F},\mathcal{C})\to(X^{\prime},\mathcal{F}^{\prime},\mathcal{C}^{\prime})$ is smooth if and only if $f\circ c\in\mathcal{C}^{\prime}$ and $f^{\prime}\circ f\in\mathcal{F}^{\prime}$ for all $f^{\prime}\in\mathcal{F}^{\prime}$, $c\in\mathcal{C}$.
This notion, introduced in [20], was later formalized in [32]. For an early comparison between diffeological and Frölicher settings, see [36]; see also [2, 37, 38, 39, 47]. In particular, [39] explains that:
Diffeological, Frölicher, and Gâteaux smoothness coincide on Fréchet spaces.
To construct a Frölicher structure on a set $X$, one can proceed as follows: given a set of generating functions $\mathcal{F}_{g}\subset\mathrm{Map}(X,\mathbb{R})$, define
$\mathcal{C}=\{c:\mathbb{R}\to X\mid f\circ c\in C^{\infty}(\mathbb{R},\mathbb{R})\ \forall f\in\mathcal{F}_{g}\},\quad\mathcal{F}=\{f:X\to\mathbb{R}\mid f\circ c\in C^{\infty}(\mathbb{R},\mathbb{R})\ \forall c\in\mathcal{C}\}.$
We say that $\mathcal{F}_{g}$ generates the Frölicher structure $(X,\mathcal{F},\mathcal{C})$.
This structure induces a natural topology on $X$, namely the pullback of the standard topology on $\mathbb{R}$ via the maps in $\mathcal{F}$.
Alternatively, one can begin with a set of generating contours $\mathcal{C}_{g}$, and define:
$\mathcal{F}=\{f:X\to\mathbb{R}\mid f\circ c\in C^{\infty}(\mathbb{R},\mathbb{R})\ \forall c\in\mathcal{C}_{g}\},\quad\mathcal{C}=\{c:\mathbb{R}\to X\mid f\circ c\in C^{\infty}(\mathbb{R},\mathbb{R})\ \forall f\in\mathcal{F}\}.$
If $X$ is a finite-dimensional smooth manifold, taking $\mathcal{F}=C^{\infty}(X,\mathbb{R})$ and $\mathcal{C}=C^{\infty}(\mathbb{R},X)$ yields a Frölicher structure whose topology matches the standard manifold topology [32]. Moreover, one can define a diffeology on a Frölicher space by:
(1) $\mathcal{P}_{\infty}(\mathcal{F}):=\coprod_{p\in\mathbb{N}^{*}}\left\{f:D(f)\subset\mathbb{R}^{p}\to X\ \middle|\ f\text{ is smooth and }\mathcal{F}\circ f\subset C^{\infty}(D(f),\mathbb{R})\right\}.$
This is known as the nebulae diffeology (cf. [31]).
Proposition 5 ([36]).
Let $(X,\mathcal{F},\mathcal{C})$ and $(X^{\prime},\mathcal{F}^{\prime},\mathcal{C}^{\prime})$ be Frölicher spaces. A map $f:X\to X^{\prime}$ is smooth in the Frölicher sense if and only if it is smooth for the associated diffeologies $\mathcal{P}_{\infty}(\mathcal{F})$ and $\mathcal{P}_{\infty}(\mathcal{F}^{\prime})$.
Hence, we have the chain of implications:
[…]
These implications can be refined and are further analyzed in [47, 36].
Remark 6.
The set of contours $\mathcal{C}$ of a Frölicher space $(X,\mathcal{F},\mathcal{C})$ does not define a diffeology by itself, because a diffeology requires closure under domain restriction. However, one can construct a minimal diffeology $\mathcal{P}_{1}(\mathcal{F})$ consisting of plots locally of the form $c\circ g$, with $c\in\mathcal{C}$ and $g$ a smooth map. Then, Proposition 5 remains valid if we replace $\mathcal{P}_{\infty}$ by $\mathcal{P}_{1}$. This is based on Boman’s theorem [32, p. 26], and detailed in [37, 36, 47].
Definition 7 ([47]).
A diffeology $\mathcal{P}_{\infty}$ defined by a set of functions $\mathcal{F}$ along the lines of (1) is called reflexive. Such diffeologies are those generated by the smooth functions of a Frölicher structure.
We now present additional examples of diffeologies.
Example 8.
Let $k\in\mathbb{N}$ and let $(X,\mathcal{F},\mathcal{C})$ be a Frölicher space. The $C^{k}$-diffeology $\mathcal{P}_{(k)}$ is defined by:
$\forall d\in\mathbb{N}^{*},\ \forall O\subset\mathbb{R}^{d}\text{ open,}\quad\mathcal{P}_{(k)}(O)=\left\{p:O\to X\mid\mathcal{F}\circ p\subset C^{k}(O,\mathbb{R})\right\},$
and we set
$\mathcal{P}_{(k)}=\bigcup_{d\in\mathbb{N}^{*}}\bigcup_{\text{open }O\subset\mathbb{R}^{d}}\mathcal{P}_{(k)}(O).$
This construction generalizes the finite-dimensional case treated in [31].
Proposition 9 ([31, 46]).
Let $(X,\mathcal{P})$ and $(X^{\prime},\mathcal{P}^{\prime})$ be diffeological spaces. The product diffeology on $X\times X^{\prime}$ is defined as the coarsest diffeology for which the projections are smooth. It consists of maps $g:O\to X\times X^{\prime}$ that decompose as $g=(f,f^{\prime})$ with $f\in\mathcal{P}$, $f^{\prime}\in\mathcal{P}^{\prime}$. This definition extends to arbitrary (even uncountable) products.
Proposition 10.
Let $(X,\mathcal{F},\mathcal{C})$ and $(X^{\prime},\mathcal{F}^{\prime},\mathcal{C}^{\prime})$ be Frölicher spaces, with associated diffeologies $\mathcal{P}$ and $\mathcal{P}^{\prime}$. Then $X\times X^{\prime}$ admits a natural Frölicher structure whose set of contours is $\mathcal{C}\times\mathcal{C}^{\prime}$, and whose associated diffeology is the product $\mathcal{P}\times\mathcal{P}^{\prime}$.
This construction also extends to infinite products.
Let $(X,\mathcal{P})$ be a diffeological space and $X^{\prime}$ a set. Given a map $f:X\to X^{\prime}$, the pushforward diffeology $f_{*}(\mathcal{P})$ is the coarsest diffeology on $X^{\prime}$ making $f$ smooth. Conversely, if $(X^{\prime},\mathcal{P}^{\prime})$ is diffeological and $f:X\to X^{\prime}$, the pullback diffeology $f^{*}(\mathcal{P}^{\prime})$ is the finest diffeology on $X$ for which $f$ is smooth.
Proposition 11 ([31, 46]).
Let $(X,\mathcal{P})$ be a diffeological space and $\sim$ an equivalence relation. Then the quotient $X/\sim$ inherits a diffeology $\mathcal{P}/\sim$ defined as the pushforward of $\mathcal{P}$ via the projection $X\to X/\sim$. Locally, plots are of the form $\pi\circ p$, with $p\in\mathcal{P}$.
Let $X_{0}\subset X$ be a subset.
  • If $X$ is diffeological with $\mathcal{P}$, the subset diffeology on $X_{0}$ is the restriction $\mathcal{P}_{0}=\{p\in\mathcal{P}\mid\mathrm{Im}(p)\subset X_{0}\}$.
  • If $(X,\mathcal{F},\mathcal{C})$ is Frölicher, we define the induced structure on $X_{0}$ using the restrictions of maps in $\mathcal{F}$ and contours in $\mathcal{C}$ with image in $X_{0}$.
Example 12.
Let $X,Y$ be sets and consider the diffeological space $Y^{X}$ with the functional product diffeology, defined as the largest diffeology making all evaluation maps $\mathrm{ev}_{x}:Y^{X}\to Y$ smooth. Any subspace of $Y^{X}$ admitting this property inherits a subdiffeology of this functional product diffeology.
Proposition 13 ([31]).
Let $X,Y,Z$ be diffeological spaces. Then the canonical identifications
$C^{\infty}(X\times Y,Z)\simeq C^{\infty}(X,C^{\infty}(Y,Z))\simeq C^{\infty}(Y,C^{\infty}(X,Z))$
hold, where all function spaces are equipped with the functional diffeology.
1.2. Groups, vector spaces, algebra and fiber bundles in diffeologies
Given an algebraic structure, one can naturally define a compatible diffeological or Frölicher structure. For instance, see [34, 37], and more specifically [31, pp. 66–68] for the case of vector spaces.
If $\mathbb{R}$ is equipped with its standard diffeology (or Frölicher structure), then an $\mathbb{R}$-vector space $V$ equipped with a diffeology (or Frölicher structure) is called a diffeological (resp. Frölicher) vector space if vector addition and scalar multiplication are smooth maps.
Definition 14.
Let $G$ be a group endowed with a diffeology (resp. a Frölicher structure). Then $G$ is called a diffeological (resp. Frölicher) group if both the multiplication map $G\times G\to G$ and the inversion map $G\to G$ are smooth.
This notion applies similarly to rings, modules, algebras, and other algebraic structures where compatibility with the diffeological or Frölicher framework requires the corresponding algebraic operations to be smooth.
Moreover, one can define fiber bundles in the diffeological context. A diffeological fiber bundle is a surjective smooth map $\pi:E\to B$ between diffeological spaces such that locally (in the diffeological sense), $E$ is diffeomorphic to $B\times F$, with $F$ the typical fiber, and the transition functions are smooth in the functional diffeology sense.
Many classical constructions—direct sums, duals, subspaces—have analogs in the diffeological category, but must be handled carefully since the lack of local triviality (in the usual topological sense) may affect their behavior. In particular, when dealing wit tangent and cotangent spaces often belong to extended classes called pseudo-bundles (see e.g. [42]), namely,
  • the cotangent space of a diffeological space is defined in a widely accepted way in the literature; but it appears that there is no typical fiber. The fibers are vector spaces, but they can be no-isomorphic. For example, the cross has a cotangent bundle with fibers of dimension 1 or 2. This is called a vector pseudo bundle
  • the notion of tangent bundle is even more difficult to study because many non-equivalent versions are actually proposed, and each of them has its own motivations.
For an overview on tangent spaces, cotangent spaces, vector pseudo bundles, we refer to [29].
2. Linearizations from Finite-Dimensional Manifolds to Diffeological Spaces and Infinite-Dimensional Settings
2.1. Linearizations of a Manifold
Bokobza-Haggiag introduces the notion of linearization in [7], aiming to extend local analytical constructions—such as differential operators—into a global structure on a smooth manifold $M$. Let us first recall the classical concept:
Definition 15.
A local linearization at a point $x\in M$ is a diffeomorphism
$\psi_{x}:U\subset M\longrightarrow V\subset\mathbb{R}^{n},$
such that $\psi_{x}(x)=0$ and $d\psi_{x}|_{x}:T_{x}M\to\mathbb{R}^{n}$ is the identity (under a fixed identification).
The paper [7] generalizes this to a global notion:
Definition 16 ([7]).
A (global) linearization is a smooth map $\nu:X\times X\rightarrow TX$ such that:
  1. (1)
    Let $p_{1}:X\times X\rightarrow X$ denote the projection onto the first component, and let $\pi:TX\rightarrow X$ denote the tangent bundle projection. Then the following diagram commutes:
[…]
  2. (2)
    For all $x\in X$, one has $\nu(x,x)=0$.
  3. (3)
    For all $x\in X$, the differential $D\nu(x,\cdot)|_{y=x}:T_{x}X\to T_{0}(T_{x}X)\simeq T_{x}X$ is the identity.
Using classical geometric techniques, one can derive the following properties:
  1. (1)
    By the inverse function theorem, a linearization defines a local linearization at each point $x\in X$.
  2. (2)
    There exists a neighborhood $\Omega$ of the diagonal $\Delta(X)\subset X\times X$ such that $\nu|_{\Omega}$ is a diffeomorphism onto a neighborhood of the zero section of $TX$.
  3. (3)
    The map
    $l_{x}^{y}:\xi\in T^{*}X\mapsto d_{y}\langle\nu(x,y),\xi\rangle$
    is a smooth linear map from $T^{*}_{y}X$ to $T^{*}_{x}X$.
  4. (4)
    Every manifold $X$ admits a linearization.
  5. (5)
    A manifold $X$ admits an atlas with affine transition functions if and only if there exists a linearization $\mu$ of $X$ such that, for all $(x,y,z)\in X^{3}$ with $(x,y),(y,z),(x,z)\in\Omega$, one has
    $l_{x}^{z}=l_{y}^{z}\circ l_{x}^{y}.$
We present an equivalent formulation of Definition 16 that will be helpful for further generalization.
Proposition 17.
The following condition is equivalent to Definition 16.
Define $k:T^{*}X\times X\to\mathbb{R}$ by
$k(\xi,y):=\langle\nu(x,y),\xi\rangle,\quad\text{with }x=\pi(\xi).$
Then $\nu$ is a global linearization if and only if:
  1. (1)
    $k(\xi,x)=0$ for all $\xi\in T^{*}X$ (where $x=\pi(\xi)$);
  2. (2)
    The differential $d_{y}k(\xi,y)|_{y=x}:T^{*}_{x}X\to T^{*}_{x}X$ is the identity.
2.2. Exponential Map and Linearization on a Compact Riemannian Manifold
Let $(M,g)$ be a compact, connected, orientable, boundaryless Riemannian manifold. By the Hopf–Rinow theorem, $M$ is geodesically complete, so the exponential map
$\exp_{p}:T_{p}M\to M$
is defined on the entire tangent space $T_{p}M$ for every $p\in M$.
2.2.1. Properties of the exponential map
We follow classical references such as [4, 21]. Let $v\in T_{p}M$ and consider the geodesic $\gamma(t)=\exp_{p}(tv)$. Then $d\exp_{p}|_{v}$ satisfies:
  • Identity at the origin:
    $d\exp_{p}|_{0}=\mathrm{id}_{T_{p}M}.$
  • Critical and regular points: A vector $v\in T_{p}M$ is a regular point if $d\exp_{p}|_{v}$ is invertible. Otherwise, $v$ is critical, and $\exp_{p}(v)$ is a conjugate point along the geodesic.
  • Local diffeomorphism: There exists $\varepsilon>0$ such that $\exp_{p}$ is a diffeomorphism from $B(0,\varepsilon)\subset T_{p}M$ onto its image.
  • Injectivity radius: The injectivity radius at $p$ is the supremum of $r>0$ such that $\exp_{p}$ is a diffeomorphism on $B(0,r)$. For compact $M$, this radius is uniform and strictly positive.
  • Uniform boundedness: If the Riemann curvature is uniformly bounded, then the norm of $d\exp_{p}$ can be estimated via comparison theorems (e.g., Rauch’s). There exists a uniform radius $r>0$ such that $\exp_{p}$ is a diffeomorphism on $B(0,r)\subset T_{p}M$.
2.2.2. The Exponential Map Defines a Linearization in the Sense of Bokobza-Haggiag
Let $M$ be as above. Since the injectivity radius $r>0$ is uniform, the exponential map defines a diffeomorphism:
$\exp:B_{r}=\left\{v\in TM\ \middle|\ \langle v,v\rangle<r^{2}\right\}\to V_{r}\subset M\times M,$
where $V_{r}$ is a neighborhood of the diagonal $\Delta(M)$ in $M\times M$.
Define $\delta:TM\to V_{r}$ by
$\delta(v):=\left(x,\exp_{x}\left(\frac{r}{\sqrt{1+\langle v,v\rangle}}v\right)\right),$
where $v\in T_{x}M$.
Then:
  • $\delta\in C^{\infty}(TM,M\times M)$;
  • For any linearization $\nu$ of $M$ in the sense of Bokobza-Haggiag, we have:
    $D_{0}\left(\left(\frac{1}{r}\nu\right)\circ\delta\right)=\mathrm{Id}_{T_{x}X}=D_{x}\left(\delta\circ\left(\frac{1}{r}\nu\right)\right),$
    where $D_{0}$ is the derivative along the fibers of $T_{X}$ and $D_{x}$ is the derivative along the second variable of $X\times X.$
Using the Riemannian metric, we identify $TM\cong T^{*}M$. Hence, both tangent and cotangent bundles may be used interchangeably in the characterization of linearizations.
These properties will serve as a reference point when generalizing linearizations to diffeological spaces.
3. Extending Classical Optimization Methods in Linearizations
3.1. A Generalized Notion of Linearization
Definition 18.
Let $(X,\mathcal{P})$ be a diffeological space. Let $E$ be a diffeological vector pseudo-bundle over $X$ with projection $\pi_{E}$, not necessarily equipped with local trivializations or smooth partitions of unity, but such that the inclusion map $T^{*}X\subset E$ is a morphism of diffeological pseudo-bundles. A linearization of $X$ is a pair $(\nu,\delta)$ of smooth maps
$\delta:E\longrightarrow X\times X,\qquad\nu:X\times X\longrightarrow E,$
such that the following properties hold:
  1. (1)
    Denoting by $\Delta(X)$ the diagonal of $X\times X,$ $\delta^{-1}_{|T^{*}X}(\Delta(X))$ is the zero section of $T^{*}X$, i.e. $\forall\xi\in T^{*}X\subset E,\delta(\xi)\in\Delta(X)\Leftrightarrow\xi=0.$
  2. (2)
    the diagram
[…]
    commutes, i.e. the base point is preserved.
  3. (3)
    $\nu:X\times X\rightarrow E$ is smooth,
  4. (4)
    $\forall(x,y)\in X^{2},$ we have $\nu(x,y)=0\Leftrightarrow x=y$,
  5. (5)
    $\forall x\in X,$ the derivative $d_{x}\nu(x,.)\neq 0$, where $d_{x}$ is the de Rham differential, here applied to the map $(\nu(x,.):y\mapsto\nu(x,y))\in C^{\infty}(X,E_{x}).$ Therefore, $d_{x}\nu(x,.)\in\Omega^{1}(X,E_{x}).$
  6. (6)
    The diagram
[…]
    commutes,
  7. (7)
    For every $x\in X$, the induced map from $T^{*}X$ to $T^{*}X$ by $(\delta\circ\nu)(x,.)\in C^{\infty}(X,X)$ restricts to an injective map on $T^{*}_{x}X$.
We denote such a structure by the pair $(\delta,\nu)$ and say that $(E,\delta,\nu)$ is a linearization of $X$.
This definition mirrors the notation in Section 2.2.2 to highlight differences with the classical definition by Bokobza-Haggiag. Notably, condition (4) does not generally hold in the classical case when $\nu$ targets $TX$ or $T^{*}X$. The non-trivial homotopy type of $X$ is a main obstruction such a global property.
This technical obstacle is circumvented by enlarging the target space to a pseudo-bundle $E$ over $X$ containing $T^{*}X$ as a sub-pseudo-bundle. The main motivation is that $T^{*}X$ is well-defined in the diffeological setting, unlike the tangent bundle, for which several constructions exist. We now verify the existence of such linearizations in standard cases, namely, when $X$ is a (classical) smooth manifold:
Proposition 19 (Smooth linearization on a differentiable manifold).
Let $(X,g)$ be a smooth finite-dimensional Riemannian manifold. Using the musical isomorphism $v\mapsto v^{\flat}:=g_{x}(v,\cdot)$, we identify $T_{x}X$ with $T_{x}^{\ast}X$. Set
$E:=T^{\ast}X\oplus\mathbb{R},$
viewed as a smooth vector bundle over $X$ via the projection $\pi_{1}:E\to X$.
Then there exists a smooth map
$\nu:X\times X\longrightarrow E$
such that:
  1. (i)
    $\nu(x,x)=(0_{x},0)$ for all $x\in X$;
  2. (ii)
    $\nu(x,y)=(0_{x},0)$ if and only if $y=x$;
  3. (iii)
    for each $x\in X$, the differential of the first component of $\nu$ with respect to the second variable, evaluated at $(x,x)$, is the musical isomorphism
    $D_{2}\bigl(pr_{1}\circ\nu\bigr)(x,x)=(\cdot)^{\flat}:T_{x}X\longrightarrow T_{x}^{\ast}X,$
    where $pr_{1}:E\to T^{\ast}X$ denotes the first projection.
In particular, $(E,\nu)$ defines a smooth linearization of $X$ along the diagonal in the above sense.
Proof.
Step 1: A Bokobza–Haggiag type linear part in $T^{\ast}X$.
Fix a smooth Riemannian metric $g$ on $X$. For each $x\in X$, let $\exp_{x}:\mathcal{U}_{x}\subset T_{x}X\to X$ be the exponential map, defined on some neighbourhood $\mathcal{U}_{x}$ of $0_{x}$. By the tubular neighbourhood theorem for the diagonal
$\Delta:=\{(x,x):x\in X\}\subset X\times X,$
there exists an open neighbourhood $U$ of $\Delta$ in $X\times X$ and a smooth map
$\xi:U\longrightarrow TX$
such that:
  1. (1)
    $\xi(x,y)\in T_{x}X$ for all $(x,y)\in U$;
  2. (2)
    $\xi(x,x)=0_{x}$ for all $x\in X$;
  3. (3)
    $y=\exp_{x}\bigl(\xi(x,y)\bigr)$ for all $(x,y)\in U$.
In particular,
$D_{2}\xi(x,x)=\mathrm{Id}_{T_{x}X}.$
Define
$\nu_{\mathrm{BH}}:U\longrightarrow T^{\ast}X,\qquad\nu_{\mathrm{BH}}(x,y):=\xi(x,y)^{\flat}.$
Then $\nu_{\mathrm{BH}}$ is smooth on $U$, satisfies $\nu_{\mathrm{BH}}(x,x)=0_{x}$, and
$D_{2}\nu_{\mathrm{BH}}(x,x)=(\cdot)^{\flat}:T_{x}X\to T_{x}^{\ast}X.$
Step 2: A smooth function vanishing exactly on the diagonal.
Define on $U$:
$\rho_{0}(x,y):=\|\xi(x,y)\|_{g}^{2}.$
Then $\rho_{0}$ is smooth on $U$, nonnegative, and
$\rho_{0}(x,y)=0\Longleftrightarrow(x,y)\in\Delta\cap U.$
Choose an open subset $U_{0}$ such that
$\Delta\subset U_{0}\subset\overline{U_{0}}\subset U,$
and a smooth cutoff function $\chi\in C^{\infty}(X\times X,[0,1])$ with
$\chi\equiv 1\ \text{on }U_{0},\qquad\mathrm{supp}(\chi)\subset U.$
Define a global smooth map
$\rho:X\times X\longrightarrow[0,+\infty)$
by
$\rho(x,y):=\chi(x,y)\,\rho_{0}(x,y)+\bigl(1-\chi(x,y)\bigr).$
Then:
  • on $U_{0}$, $\rho=\rho_{0}$, hence $\rho(x,y)=0$ iff $(x,y)\in\Delta$;
  • outside $U$, $\chi=0$ and thus $\rho\equiv 1$.
Hence $\rho$ is smooth and
$\rho(x,y)=0\Longleftrightarrow(x,y)\in\Delta.$
Step 3: Global extension of $\nu_{\mathrm{BH}}$.
Define
$\widetilde{\nu}_{\mathrm{BH}}(x,y):=\chi(x,y)\,\nu_{\mathrm{BH}}(x,y),$
extending $\nu_{\mathrm{BH}}$ smoothly across $U$ and taking it to 0 outside $U$. Thus $\widetilde{\nu}_{\mathrm{BH}}$ is smooth on $X\times X$, and coincides with $\nu_{\mathrm{BH}}$ on $U_{0}$.
Step 4: Definition of the linearization $\nu$.
Set
$\nu(x,y):=\bigl(\widetilde{\nu}_{\mathrm{BH}}(x,y),\;\rho(x,y)\bigr)\in T_{x}^{\ast}X\oplus\mathbb{R}.$
Step 5: Verification of properties (i)–(iii).
(i) For $y=x$, we have $(x,x)\in U_{0}$. Hence $\chi(x,x)=1$, $\rho(x,x)=\|\xi(x,x)\|^{2}=0$, and $\widetilde{\nu}_{\mathrm{BH}}(x,x)=\nu_{\mathrm{BH}}(x,x)=0_{x}$. Thus
$\nu(x,x)=(0_{x},0).$
(ii) If $\nu(x,y)=(0_{x},0)$, then in particular $\rho(x,y)=0$. By construction of $\rho$, this implies $(x,y)\in\Delta$, hence $y=x$. The converse is already proved in (i). Thus
$\nu(x,y)=(0_{x},0)\quad\Longleftrightarrow\quad y=x.$
(iii) Near $(x,x)$, we are in $U_{0}$, so
$pr_{1}\circ\nu=\widetilde{\nu}_{\mathrm{BH}}=\nu_{\mathrm{BH}}.$
Therefore
$D_{2}(pr_{1}\circ\nu)(x,x)=D_{2}\nu_{\mathrm{BH}}(x,x)=(\cdot)^{\flat}:T_{x}X\to T_{x}^{\ast}X,$
which is the required linear isomorphism.
This completes the proof. ∎
Proposition 20 (Construction of $\delta$ compatible with $\nu$).
Let $(X,g)$ be a smooth finite-dimensional Riemannian manifold, and let
$E:=T^{\ast}X\oplus\mathbb{R}$
as in Proposition 19. Let $\nu:X\times X\to E$ be the smooth map constructed there, i.e.
$\nu(x,y)=\bigl(\widetilde{\nu}_{\mathrm{BH}}(x,y),\,\rho(x,y)\bigr),$
where $\widetilde{\nu}_{\mathrm{BH}}$ extends the Bokobza–Haggiag-type linearization and $\rho$ is a smooth function vanishing exactly on the diagonal $\Delta\subset X\times X$.
Then there exists a smooth map
$\delta:E\longrightarrow X\times X$
such that:
  1. (i)
    $\delta(x,0_{x},s)=(x,x)$ for all $x\in X$ and $s\in\mathbb{R}$;
  2. (ii)
    there exists an open neighbourhood $U_{0}$ of the diagonal in $X\times X$ such that, for all $(x,y)\in U_{0}$, writing
    $\nu(x,y)=\bigl(\alpha_{x,y},\,\rho(x,y)\bigr)\in T_{x}^{\ast}X\oplus\mathbb{R},$
    one has
    $\delta\bigl(x,\alpha_{x,y},\rho(x,y)\bigr)=(x,y);$
  3. (iii)
    for each $x\in X$, the vertical differential of $\delta$ at $(x,0_{x},0)$, restricted to the $T_{x}^{\ast}X$-component, induces the musical isomorphism
    $D^{\mathrm{vert}}_{\alpha}\delta(x,0_{x},0):T_{x}^{\ast}X\xrightarrow{\ \sim\ }\{0\}\oplus T_{x}X$
    via the identification $T_{(x,x)}(X\times X)\simeq T_{x}X\oplus T_{x}X$.
Consequently, $(E,\nu,\delta)$ defines a smooth linearization of $X$ in the sense that, near the diagonal, $\nu$ and $\delta$ are local inverses in the second variable and encode the tangent structure via $g$.
Proof.
We use the same geometric data as in Proposition 19. In particular, we have:
  • an open neighbourhood $U\subset X\times X$ of the diagonal $\Delta$;
  • a smooth map $\xi:U\to TX$ such that $\xi(x,y)\in T_{x}X$, $\xi(x,x)=0_{x}$, and $y=\exp_{x}(\xi(x,y))$;
  • an open subset $U_{0}$ with $\Delta\subset U_{0}\subset\overline{U_{0}}\subset U$;
  • a cutoff function $\chi\in C^{\infty}(X\times X,[0,1])$ with $\chi\equiv 1$ on $U_{0}$ and $\mathrm{supp}(\chi)\subset U$;
  • the function $\rho(x,y)=\chi(x,y)\,\|\xi(x,y)\|_{g}^{2}+(1-\chi(x,y))$;
  • the extended Bokobza–Haggiag map $\widetilde{\nu}_{\mathrm{BH}}(x,y)=\chi(x,y)\,\xi(x,y)^{\flat}$.
Step 1: Controlling the exponential domain.
For each $x\in X$, choose $r(x)>0$ such that the open ball
$B_{x}:=\{v\in T_{x}X:\|v\|_{g}<r(x)\}$
is contained in the domain of $\exp_{x}$ and $\exp_{x}:B_{x}\to\exp_{x}(B_{x})$ is a diffeomorphism onto its image. By standard arguments (using a partition of unity), we may assume that $x\mapsto r(x)$ is a smooth positive function.
Let $\psi:[0,+\infty)\to[0,1]$ be a smooth function such that
$\psi(t)=1\quad\text{for }t\leq 1,\qquad\psi(t)=0\quad\text{for }t\geq 2.$
Step 2: Definition of $\delta$.
For $(x,\alpha,s)\in E_{x}=T_{x}^{\ast}X\oplus\mathbb{R}$, set
$v:=\alpha^{\sharp}\in T_{x}X,\qquad\lambda(x,\alpha):=\psi\!\left(\frac{\|v\|_{g}}{r(x)}\right),\qquad v^{\prime}(x,\alpha):=\lambda(x,\alpha)\,v.$
Then $v^{\prime}(x,\alpha)$ is smooth in $(x,\alpha)$, and:
  • if $\|v\|_{g}\leq r(x)$, then $\lambda(x,\alpha)=1$ and $v^{\prime}=v$;
  • if $\|v\|_{g}\geq 2\,r(x)$, then $\lambda(x,\alpha)=0$ and $v^{\prime}=0$.
We now define
$\delta:E\longrightarrow X\times X,\qquad\delta(x,\alpha,s):=\bigl(x,\exp_{x}(v^{\prime}(x,\alpha))\bigr).$
Since $\exp_{x}$ is smooth in both $x$ and $v$, and $v^{\prime}$ is smooth, $\delta$ is a smooth map.
Step 3: Verification of (i).
If $\alpha=0_{x}$, then $v=0_{x}$ and hence $v^{\prime}(x,0_{x})=0_{x}$ for all $x$. Therefore
$\delta(x,0_{x},s)=(x,\exp_{x}(0_{x}))=(x,x),$
for all $s\in\mathbb{R}$. This proves (i).
Step 4: Compatibility with $\nu$ near the diagonal.
Recall that, on $U_{0}$, we have $\chi\equiv 1$, hence
$\nu(x,y)=\bigl(\nu_{\mathrm{BH}}(x,y),\rho(x,y)\bigr)=\bigl(\xi(x,y)^{\flat},\|\xi(x,y)\|_{g}^{2}\bigr).$
Fix $(x,y)\in U_{0}$ and set
$\alpha_{x,y}:=\xi(x,y)^{\flat}.$
Then
$v:=\alpha_{x,y}^{\sharp}=\xi(x,y),\qquad\|v\|_{g}=\|\xi(x,y)\|_{g}.$
For $(x,y)$ sufficiently close to the diagonal, we may assume that $\|v\|_{g}<r(x)$, so that $\lambda(x,\alpha_{x,y})=1$ and $v^{\prime}(x,\alpha_{x,y})=v$. Hence
$\delta\bigl(x,\alpha_{x,y},\rho(x,y)\bigr)=\bigl(x,\exp_{x}(v^{\prime}(x,\alpha_{x,y}))\bigr)=\bigl(x,\exp_{x}(\xi(x,y))\bigr)=(x,y).$
This establishes (ii) (possibly after shrinking $U_{0}$ if needed).
Step 5: Vertical differential at the zero section.
Consider the vertical derivative of $\delta$ at $(x,0_{x},0)$, restricted to the $T_{x}^{\ast}X$-component. For $\alpha$ small, we have $v=\alpha^{\sharp}$ small in $T_{x}X$, so that $\|v\|_{g}<r(x)$ and $\lambda(x,\alpha)=1$. Hence $v^{\prime}(x,\alpha)=v$ in a neighbourhood of $(x,0_{x})$, and
$\delta(x,\alpha,0)=\bigl(x,\exp_{x}(\alpha^{\sharp})\bigr).$
Differentiating at $\alpha=0_{x}$, we obtain
$D_{\alpha}\delta(x,0_{x},0)(\beta)=\bigl(0,\,D\exp_{x}(0_{x})(\beta^{\sharp})\bigr)=\bigl(0,\,\beta^{\sharp}\bigr),$
for all $\beta\in T_{x}^{\ast}X$, since $D\exp_{x}(0_{x})=\mathrm{Id}_{T_{x}X}$. Under the identification $T_{(x,x)}(X\times X)\simeq T_{x}X\oplus T_{x}X$, this means that $D^{\mathrm{vert}}_{\alpha}\delta(x,0_{x},0)$ maps $T_{x}^{\ast}X$ isomorphically onto the subspace $\{0\}\oplus T_{x}X$ via $\beta\mapsto(0,\beta^{\sharp})$, which proves (iii).
This shows that $\delta$ has the required properties, and that, near the diagonal, $\delta$ and $\nu$ are local inverses in the second variable with the correct linear behaviour encoded by the Riemannian metric. ∎
3.2. Distance Induced by a Linearization
We now introduce an intrinsic structure for measuring convergence on $X$ using the linearization $(\delta,\nu)$. Note that the map $\delta$ plays a central role in this construction.
Definition 21 (Distance induced by a linearization).
Let $X$ be a diffeological space equipped with a linearization $(\delta,\nu)$ and a smooth function $\mathrm{gap}:E\to\mathbb{R}$ (called the gap function) satisfying:
  • (i)
    $\mathrm{gap}(0_{x})=0$ for all $x\in X$;
  • (ii)
    $\mathrm{gap}(\nu(x,y))=0$ if and only if $x=y$;
  • (iii)
    for each fixed $x$, the map $y\mapsto\mathrm{gap}(\nu(x,y))$ is smooth.
Then the functions
$d_{X}:(x,y)\in X^{2}\mapsto\sup_{z\in X}\big|\mathrm{gap}(\nu(z,x))-\mathrm{gap}(\nu(z,y))\big|,$
$\forall x\in x,d_{E}:(u,v)\in E_{x}^{2}\mapsto d_{X}(\delta(u-v))+d_{X}(\delta(v-u)).$
define pseudo-distances on $X$ and on each fiber of $E$, respectively. When these pseudo-distances are non-degenerate, they are called the distances induced by the linearization $(\delta,\nu)$.
To simplify notation, subscripts are dropped when context permits.
Theorem 22.
Let $X$ be a diffeological space equipped with a linearization $(\delta,\nu)$ and a gap function satisfying Assumption D. Then $d_{X}$ is a distance on $X$ and $d_{E}$ is a distance on each fiber of $E$.
Proof.
We verify the axioms of a metric for $d_{X}$ and $d_{E}$:
  1. (1)
    Separating property:
    $gap(\nu(x,y))\neq 0\iff x\neq y.$
    Thus, $d_{X}(x,x)=0$, and if $x\neq y$, then choosing $z=x$ gives $d_{X}(x,y)>0$.
    Similarly, $d_{E}(u,v)==0$ if and only if $u=v$.
  2. (2)
    Symmetry: immediate from definitions.
  3. (3)
    Triangle inequality: For $x,y,z\in X$ and all $w\in X$:
    $\left|gap(\nu(w,x))-gap(\nu(w,z))\right|$ $\leq$ $\left|gap(\nu(w,x))-gap(\nu(w,y))\right|$
    $+\left|gap(\nu(w,y))-gap(\nu(w,z))\right|.$
    Taking suprema yields $d_{X}(x,z)\leq d_{X}(x,y)+d_{X}(y,z)$.
Remark 23.
Convexity of the gap function is not required for these properties.
Definition 24.
Let $\tilde{X}$ denote the metric completion of $X$ with respect to $d_{X}$.
4. Gradient and steepest descent methods… (Almost) Without gradient
Let us now consider a minimization problem of a given functional $F:X\rightarrow\mathbb{R}$ or an equation $F(u)=0$ on a diffeological space [^1] We recall that solving $F(x)=0$ is equivalent to solving the minimization problem for $|F|.$. We define the following basic notion:
Definition 25.
A method $\mathcal{M}$ is a mapping that depends on:
  • a smooth function $f\in C^{\infty}(I;\mathbb{R})$, where $I\subset\mathbb{R}$ is compact,
  • and a point $x\in I$,
and produces a point $\mathcal{M}_{f}(x)\in I$ such that:
  1. (1)
    $\forall x\in I,\quad f(\mathcal{M}_{f}(x))\leq f(x)$,
  2. (2)
    $\forall x\in I,\quad f(\mathcal{M}_{f}(x))=f(x)\iff\mathcal{M}_{f}(x)=x$,
  3. (3)
    $\mathcal{M}_{f}$ is smooth on any subinterval $J\subset I$ where $df\neq 0$.
By this definition, we intend to define a method of the type of the gradient methods, Newton methods for convex functions, and more generally any descent method with exact line search that (often approximatively at a finite number of iterations) intend to produce the minimum of a smooth function $f$ of prescribed type on a compact interval.
4.1. The Algorithm Associated to a Method $\mathcal{M}$
To provide context, we first sketch the general principle of the method. The core idea is to build a smooth curve $\gamma$ on $X$, replacing the gradient direction used in classical optimization by a path derived from the linearization.
This path is defined by the linearization map $\delta$, such that:
$\gamma(t)=\delta(t\,d_{x}F).$
Data: Initial point $x_{0}\in X$, maximal number of iterations $N_{\max}$, compact search interval $I=[0,T]\subset\mathbb{R}$.
Result: A sequence $(x_{n})_{n\geq 0}\subset X$ with non-increasing values of $F$.
for $n=0,1,\dots,N_{\max}$ do
  // Construction of a search curve from the linearization at $x_{n}$
  Construct a smooth curve $\gamma_{n}:I\to X$ with $\gamma_{n}(0)=x_{n}$ using the linearization $(E,\delta,\nu)$ at $x_{n}$;
 // Model or evaluation of $F$ along the curve
  Define the real-valued function
$\varphi_{n}(t):=F(\gamma_{n}(t)),\qquad t\in I.$
Optionally, introduce a method $\mathcal{M}_{F}(\gamma_{n}):I\to\mathbb{R}$ approximating $\varphi_{n}$, and use it only to initialise the line search;
 // Exact line search on a compact interval
  Choose $t_{n}\in I$ such that
$\varphi_{n}(t_{n})\leq\varphi_{n}(t)\quad\text{for all }t\in I\quad\text{(exact minimizer on $I$).}$
 // Update of the iterate
  Set
$x_{n+1}:=\gamma_{n}(t_{n}).$
 // Monotonicity of the objective
  Then $F(x_{n+1})=\varphi_{n}(t_{n})\leq\varphi_{n}(0)=F(x_{n})$;
Algorithm 1 Diffeological line-search method induced by a linearization
Remark 26.
Our algorithm is an adaptation of the algorithm presented in [26, 27, 28], where the authors use a retraction map from $TX$ to $X$ in place of $\delta$, and replace the gradient by the differential dfdf in $T^{*}X$.
4.2. Classical A Priori Convergence Results to Weak Solutions
Definition 27 (Algorithmic stationarity).
Let $X$ be a diffeological space and $F:X\to\mathbb{R}$ a function. Assume that to each $x\in X$ is associated a smooth search curve $\gamma_{x}:I\to X$ on a fixed compact interval $I=[0,T]$, with $\gamma_{x}(0)=x$. We say that $x$ is algorithmically stationary (with respect to $(\gamma_{x})_{x}$ and $F$) if for every such curve one has
$F(\gamma_{x}(0))\leq F(\gamma_{x}(t))\quad\text{for all }t\in I.$
Equivalently, $t=0$ is a minimizer of $t\mapsto F(\gamma_{x}(t))$ on $I$.
Theorem 28 (Limit behaviour of the line-search sequence).
Assume:
  1. (H1)
    $X$ is continuously embedded into a complete metric vector space $(Y,\|\cdot\|)$, and the closure $\overline{X}^{Y}$ of $X$ in $Y$ is compact.
  2. (H2)
    $F:X\to\mathbb{R}$ admits a continuous extension (still denoted $F$) to $\overline{X}^{Y}$.
  3. (H3)
    There exists a smooth map
    $\Gamma:X\times I\longrightarrow X,\qquad(x,t)\mapsto\gamma_{x}(t),$
    where $I=[0,T]$ is a fixed compact interval, such that for each $x\in X$, the curve $t\mapsto\gamma_{x}(t)$ is the search curve used in Algorithm 1, and $\gamma_{x}(0)=x$.
  4. (H4)
    Algorithm 1 performs an exact line search on $I$, i.e. for each $n$ it chooses $t_{n}\in I$ such that
    $F(\gamma_{x_{n}}(t_{n}))\leq F(\gamma_{x_{n}}(t))\quad\text{for all }t\in I,$
    and sets $x_{n+1}:=\gamma_{x_{n}}(t_{n})$.
Let $(x_{n})_{n\geq 0}$ be the sequence generated by Algorithm 1 from an initial point $x_{0}\in X$. Then:
  1. (i)
    The sequence $(F(x_{n}))_{n\geq 0}$ is non-increasing and converges to a limit $L\in(-\infty,F(x_{0})]$.
  2. (ii)
    The set $S$ of cluster points of $(x_{n})$ in $\overline{X}^{Y}$ is non-empty and compact. Moreover,
    $F(x)=L\quad\text{for every }x\in S.$
    In particular, $F$ is constant on $S$.
  3. (iii)
    Every cluster point $x^{\star}\in S\cap X$ is algorithmically stationary in the sense of Definition 27 with respect to $F$ and the family of curves $\gamma_{x}$.
Proof.
We proceed in three steps.
Step 1: Monotonicity and convergence of $(F(x_{n}))$. By construction of Algorithm 1 and assumption (H4), we have for each $n\geq 0$:
$F(x_{n+1})=F(\gamma_{x_{n}}(t_{n}))\leq F(\gamma_{x_{n}}(0))=F(x_{n}).$
Thus $(F(x_{n}))_{n\geq 0}$ is a non-increasing sequence in $\mathbb{R}$. By (H1) and (H2), $F$ is continuous on the compact set $\overline{X}^{Y}$, hence bounded below. Therefore $(F(x_{n}))$ admits a finite limit $L\in(-\infty,F(x_{0})]$:
$F(x_{n})\downarrow L\quad\text{as }n\to\infty.$
Step 2: Cluster points and constancy of $F$ on $S$. By (H1), the sequence $(x_{n})$ takes values in $X\subset\overline{X}^{Y}$, whose closure is compact. Hence $(x_{n})$ admits at least one cluster point in $\overline{X}^{Y}$; let $S$ denote the set of all such cluster points. Then $S$ is non-empty and compact (as a closed subset of a compact space).
Let $x^{\star}\in S$. By definition, there exists a subsequence $(x_{n_{k}})$ such that $x_{n_{k}}\to x^{\star}$ in $Y$ as $k\to\infty$. By continuity of $F$ on $\overline{X}^{Y}$ (assumption (H2)), we have
$F(x_{n_{k}})\longrightarrow F(x^{\star}).$
On the other hand, by Step 1,
$F(x_{n_{k}})\longrightarrow L,$
since $(F(x_{n}))$ converges and $(n_{k})$ is an increasing sequence of indices. Thus $F(x^{\star})=L$.
As $x^{\star}\in S$ was arbitrary, we deduce that
$F(x)=L\quad\text{for all }x\in S,$
so $F$ is constant on $S$. This proves (ii).
Step 3: Algorithmic stationarity of cluster points in $X$. Let $x^{\star}\in S\cap X$ be a cluster point which belongs to $X$. By definition of $S$, there exists a subsequence $(x_{n_{k}})$ such that
$x_{n_{k}}\longrightarrow x^{\star}\quad\text{in }Y\text{ as }k\to\infty.$
Fix the corresponding search curve $\gamma_{x^{\star}}:I\to X$ given by (H3), with $\gamma_{x^{\star}}(0)=x^{\star}$. Assume, for contradiction, that $x^{\star}$ is not algorithmically stationary. Then, by Definition 27, there exists $t_{0}\in I$ such that
$F(\gamma_{x^{\star}}(t_{0}))<F(\gamma_{x^{\star}}(0))=F(x^{\star}).$
Set
$\eta:=\frac{1}{2}\Bigl(F(x^{\star})-F(\gamma_{x^{\star}}(t_{0}))\Bigr)>0.$
By (H3), the map $\Gamma:X\times I\to X$ is smooth, hence continuous in $x$ and $t$. By (H2), $F$ is continuous in $X$ (and in $\overline{X}^{Y}$). Therefore the composite
$(x,t)\longmapsto F(\Gamma(x,t))=F(\gamma_{x}(t))$
is continuous on $X\times I$.
By continuity at $(x^{\star},t_{0})$ and $(x^{\star},0)$, there exists $\delta>0$ such that for all $x\in X$ with $\|x-x^{\star}\|<\delta$ (in the metric of $Y$) we have
$\bigl|F(\gamma_{x}(t_{0}))-F(\gamma_{x^{\star}}(t_{0}))\bigr|$ $<\eta,$
$\bigl|F(\gamma_{x}(0))-F(\gamma_{x^{\star}}(0))\bigr|$ $=\bigl|F(x)-F(x^{\star})\bigr|<\eta.$
Since $x_{n_{k}}\to x^{\star}$, there exists $K$ such that for all $k\geq K$, $\|x_{n_{k}}-x^{\star}\|<\delta$. For such $k$, we have in particular:
$F(\gamma_{x_{n_{k}}}(t_{0}))<F(\gamma_{x^{\star}}(t_{0}))+\eta=F(x^{\star})-\eta<F(x_{n_{k}})-\eta+\eta=F(x_{n_{k}}),$
where we used $F(\gamma_{x^{\star}}(t_{0}))=F(x^{\star})-2\eta$ and $|F(x_{n_{k}})-F(x^{\star})|<\eta$.
Now, by exact line search (assumption (H4)), we choose $t_{n_{k}}\in I$ with
$F(\gamma_{x_{n_{k}}}(t_{n_{k}}))\leq F(\gamma_{x_{n_{k}}}(t))\quad\text{for all }t\in I.$
In particular,
$F(x_{n_{k}+1})=F(\gamma_{x_{n_{k}}}(t_{n_{k}}))\leq F(\gamma_{x_{n_{k}}}(t_{0}))<F(x_{n_{k}}).$
By Step 1, $(F(x_{n}))$ converges to $L$. Passing to the limit $k\to\infty$ in the inequality $F(x_{n_{k}+1})<F(x_{n_{k}})$ and using continuity, we obtain
$L\leq\lim_{k\to\infty}F(x_{n_{k}+1})\leq\lim_{k\to\infty}F(\gamma_{x_{n_{k}}}(t_{0}))=F(\gamma_{x^{\star}}(t_{0}))<F(x^{\star})=L,$
which is a contradiction.
Hence our assumption was false, and $x^{\star}$ must satisfy
$F(\gamma_{x^{\star}}(0))\leq F(\gamma_{x^{\star}}(t))\quad\text{for all }t\in I,$
i.e. $x^{\star}$ is algorithmically stationary in the sense of Definition 27. This proves (iii) and completes the proof. ∎
5. Application to Mapping Spaces with Low Regularity
The notion of manifolds of smooth maps used here follows the framework in [17]; see also [32], [19]. Let $E$ be a smooth vector bundle over a compact, connected, boundaryless manifold $M$ of dimension $m$. Denote by $C^{\infty}(M,E)$ the space of smooth sections $\sigma$ of $E$.
Let $\mathbb{K}$$\mathbb{R}$ or $\mathbb{C}$, and let $N$ be a smooth Riemannian manifold without boundary. Then:
- $C^{\infty}(M,N)$ denotes the set of smooth maps $f:M\rightarrow N$,
- $C^{\infty}(M,G)$ denotes the space of smooth maps from $M$ to a Lie group $G$.
The space $C^{\infty}(M,N)$ is a smooth Fréchet manifold, and $C^{\infty}(M,G)$ is a smooth Fréchet Lie group.
The connected components of $C^{\infty}(M,N)$ correspond to the homotopy classes of based maps $f:M\rightarrow N$. Note that two such components need not be modeled on the same topological vector space. However, if $f$ and $g$ lie in the same path-connected component, then the bundles $f^{*}TN$ and $g^{*}TN$ are isomorphic. The atlas structure of $C^{\infty}(M,N)$ can be found e.g. in [35].
5.1. Spaces of Mappings with Low Regularity
In the following, we consider a complex algebra $\mathcal{A}$ equipped with the Hermitian matrix product $(A,B)\mapsto\operatorname{tr}(AB^{*})$. When unambiguous, we denote the corresponding norm by $\|\cdot\|$ or $\|\cdot\|_{\mathcal{A}}$. Defining Sobolev classes of functions between two manifolds, and studying the properties of the space of functions, is a long-standing problem, coming from the investigations on the structures of smooth infinite dimensional manifolds [17] and rapidly raised in various contexts, see e.g. [5, 6, 9, 14, 18, 44]. Depending on the definitions, the space of smooth mappings is dense, or not, in these function spaces with low regularity. For our purpose, we follow and generalize [40, 29] and we choose to follow a definition based on density arguments:
Definition 29.
Let $(s,p)\in\mathbb{R}\times(1,+\infty)$, let $M$ be a compact Riemannian manifold without boundary, and let $N$ be a Riemannian submanifold embedded in $\mathcal{M}$. We define $W^{s,p}(M,\mathcal{A})$ as the completion of $C^{\infty}(M,\mathcal{A})$ for the norm
[…]
where $\Delta$ is the positive Laplacian. Then, the space $W^{s,p}(M,N)$ is defined as the completion of $C^{\infty}(M,N)$ in $W^{s,p}(M,\mathcal{A})$ for the strong topology.
We emphasize that the classical condition $s<\dim(M)/p$ (necessary for the existence of an atlas [17]) is not assumed here. When $s\leq\dim(M)/p$, this results in a weaker differential structure on $W^{s,p}(M;\mathcal{A})$, as detailed in [40].
Proposition 30.
The space $W^{s,p}(M,N)$ is a diffeological space equipped with the subset diffeology inherited from $W^{s,p}(M,\mathcal{A})$.
This completion coincides with the $\overline{X}$ construction used in earlier sections.
Remark 31 (Completions of $C^{\infty}(M,N)$ in Sobolev-type spaces).
The choices made in Definition 29 may raise many questions and objections due to various considerations on completion, topology and even intuition on function spaces of low regularity, as already mentioned. In particular, a superficial reading may consider this definition as misleading or false considering [5, 6]. For these reasons, we feel the need to precise the few following aspects. Let $M$ and $N$ be compact smooth manifolds and $1\leq p<\infty$.
  • (a)
    For $s\in\mathbb{R}_{+}^{*}$, the Sobolev-type space $W^{s,p}(M,N)$ is classically defined as follows for $p$ large enough: $W^{s,p}(M,N)$ is the completion of $C^{\infty}(M,N)$ with respect to the metric induced by a fixed embedding $N\hookrightarrow\mathbb{R}^{k}$ and the standard Sobolev norm on $W^{s,p}(M,\mathbb{R}^{k})$, which perfectly fits with pour definition.
  • (b)
    If $s=0$, then $W^{0,p}(M,N)=L^{p}(M,N)$,
  • (c)
    If $s<0$, one can make an objection, because $W^{s,p}(M,\mathcal{A})$ is defined as the dual space $(W^{-s,p^{\prime}}(M,\mathcal{A}))^{\prime}$ and this suggests that the “good” topology for completionis the weak one whenever this makes sense, in particular, when $N$ is a compact subset of a Banach space, this dual coincides with the completion of $C^{\infty}(M,N)$ in the weak topology.
However, these are features that appear with choices in the definitions of $W^{s,p}(M,\mathcal{A})$ different from Definition 29. However, we can anyway conclude that the resulting space $W^{s,p}(M,N),$ following any of the definitions existing in the literature, inherits, as a subspace of a locally convex topological vector space, a natural diffeological structure as the subset diffeology induced from $W^{s,p}(M,\mathcal{A})$, where $\mathcal{A}$ is a linear ambient space containing $N$.
5.2. From Linearizations on $N$ to Linearizations on $C^{\infty}(M,N)$ and the Equation $F(u)=0$ in $W^{s,p}(M,N)$
Since $N$ is compact, it admits a linearization $(\delta,\nu)$ between $TN\times\mathbb{R}$ and $N\times N$.
Definition 32.
Let $f\in C^{\infty}(M,N)$. Then:
  • For all $g\in C^{\infty}(M,N)$, define $\nu(f,g):x\mapsto\nu(f(x),g(x))\in f^{*}(T^{*}N)\times\mathbb{R}$.
  • For any section $s$ of $f^{*}(T^{*}N)\times\mathbb{R}$, define $\delta(s)=(f,g)$ such that for all $x\in M$, $(f(x),g(x))=\delta(s(x))$.
Proposition 33.
The maps
$\delta\in C^{\infty}(C^{\infty}(M,TN\times\mathbb{R}),C^{\infty}(M,N)\times C^{\infty}(M,N)),\quad\nu\in C^{\infty}(C^{\infty}(M,N)\times C^{\infty}(M,N),C^{\infty}(M,TN\times\mathbb{R}))$
define a linearization of $C^{\infty}(M,N)$.
We now consider a functional equation:
$F(u)=0,\quad\text{with }F:D\subset W^{s,p}(M,N)\rightarrow\mathbb{R},$
where $D\cap C^{\infty}(M,N)$ is dense in $D$, and $D$ is closed in $W^{s,p}(M,N)$. We assume that:
Assumption R. The linearization $(\delta,\nu)$ on $C^{\infty}(M,N)$ restricts to a linearization on the subset $D\cap C^{\infty}(M,N)$ equipped with the subset diffeology.
This assumption is non-trivial in general. For instance, if $D\cap C^{\infty}(M,N)$ inherits the discrete diffeology, then $T^{*}C^{\infty}(M,N)=C^{\infty}(M,N)$ and the cotangent space has trivial fibers. This highlights the potential limitations of such restrictions, and suggests that constructing a dedicated linearization directly on $D\cap C^{\infty}(M,N)$ may be more effective for algorithmic implementations.
In what follows, we restrict ourselves to the case $D=W^{s,p}(M,N)$ and $D\cap C^{\infty}(M,N)=C^{\infty}(M,N)$.
Theorem 34 (Structural properties of $W^{s,p}(M,N)$).
Let $M$ be a compact smooth manifold of dimension $m$, and let $N\subset A$ be a compact embedded submanifold of a finite-dimensional vector space $A\simeq\mathbb{R}^{k}$. Fix $1\leq p<\infty$ and $s\in\mathbb{R}$.
  1. (i)
    For $s\geq 0$, we define $W^{s,p}(M,A)$ as the usual Sobolev–Slobodeckij (or Bessel potential) space ([Adams–Fournier, Sobolev Spaces, Chap. 7]). For $s<0$, we set
    $W^{s,p}(M,A):=\bigl(W^{-s,p^{\prime}}_{0}(M,A)\bigr)^{\ast},$
    where $p^{\prime}$ is the Hölder conjugate of $p$ and $W^{-s,p^{\prime}}_{0}(M,A)$ is the closure of $C^{\infty}_{0}(M,A)$ in $W^{-s,p^{\prime}}(M,A)$ (cf. e.g. Adams–Fournier, Thm. 7.48).
  2. (ii)
    If $s<0$, the inclusion
    $C^{0}(M,A)\hookrightarrow W^{s,p}(M,A)$
    is continuous. More precisely, since $M$ is compact, every continuous $A$-valued map is in $L^{p}(M,A)$, and the canonical injection $L^{p}(M,A)\hookrightarrow W^{s,p}(M,A)$ (viewing $L^{p}$ as a subspace of $W^{s,p}$ via distributions) is continuous; hence the composition $C^{0}(M,A)\hookrightarrow L^{p}(M,A)\hookrightarrow W^{s,p}(M,A)$ is continuous.
  3. (iii)
    The space $C^{\infty}(M,N)\subset C^{0}(M,N)$ is bounded in $C^{0}(M,N)$: indeed, since $N$ is compact, there exists $R>0$ such that $\|u(x)\|_{A}\leq R$ for all $u\in C^{\infty}(M,N)$ and all $x\in M$, so
    $\|u\|_{C^{0}(M,A)}\leq R\quad\text{for all }u\in C^{\infty}(M,N).$
    In particular, for $s<0$, the inclusion $C^{\infty}(M,N)\hookrightarrow W^{s,p}(M,A)$ is bounded by (ii).
  4. (iv)
    Define $W^{s,p}(M,N)$ as the closure of $C^{\infty}(M,N)$ in $W^{s,p}(M,A)$:
    $W^{s,p}(M,N):=\overline{C^{\infty}(M,N)}^{\,W^{s,p}(M,A)}.$
    Then $W^{s,p}(M,N)$ is a closed subset of $W^{s,p}(M,A)$ and the canonical inclusion
    $W^{s,p}(M,N)\hookrightarrow W^{s,p}(M,A)$
    is continuous and maps bounded subsets of $W^{s,p}(M,N)$ into bounded subsets of $W^{s,p}(M,A)$.
No general relative compactness statement for $W^{s,p}(M,N)$ in $W^{s,p}(M,A)$ is claimed here; in our applications we instead assume directly that the relevant sublevel sets of the functional are relatively compact in $W^{s,p}(M,A)$.
Proof.
Item (i) is the standard definition of Sobolev spaces of positive and negative order; see e.g. Adams–Fournier, Sobolev Spaces, Chap. 7, in particular Thm. 7.48 for the duality characterization when $s<0$.
For (ii), since $M$ is compact, the inclusion $C^{0}(M,A)\hookrightarrow L^{p}(M,A)$ is continuous. The identification of $L^{p}(M,A)$ as a subspace of $W^{s,p}(M,A)$ for $s<0$ is again standard in the theory of negative-order Sobolev spaces, and the corresponding inclusion is continuous; composing both inclusions yields the claim.
For (iii), the image of $M$ under any $u\in C^{\infty}(M,N)$ is contained in the compact set $N$, hence the $C^{0}$-norm of $u$ is uniformly bounded in terms of $\sup_{y\in N}\|y\|_{A}$. Together with (ii), this yields boundedness in $W^{s,p}(M,A)$ when $s<0$.
Item (iv) follows from the definition of $W^{s,p}(M,N)$ as a closure in the Banach space $W^{s,p}(M,A)$: closedness and continuity of the inclusion are immediate. The last sentence is included to emphasize that, for the purposes of our convergence results, we do not rely on any general compactness property of $W^{s,p}(M,N)$ in $W^{s,p}(M,A)$, but instead assume compactness at the level of sublevel sets. ∎
6. Outlook: fixed-point perspectives
The present work develops a diffeological framework for line-search methods based solely on evaluations of the objective function along search curves induced by a linearization. The resulting convergence theory establishes monotonicity of the objective values and algorithmic stationarity of cluster points under very mild assumptions.
A natural direction for future research is to reinterpret the update rule $x_{n+1}=\gamma_{x_{n}}(t_{n})$ as a fixed-point iteration $x_{n+1}=T(x_{n})$ generated by the linearization; here $T$ depends nonlinearly on both the choice of the search curve and on the outcome of the line search. Such interpretations arise in classical optimization when the search direction is derived from a gradient or Newton-type model.
In the present diffeological context, establishing a genuine fixed-point theory for the map $T$ would require additional hypotheses ensuring, for instance, compactness or contractivity in a complete metric model. These assumptions are well outside the scope of the present article and are not needed for our convergence results. Developing such a theory could nevertheless be useful in applications to nonlinear PDEs or variational problems where a diffeological linearization carries geometric information; we leave this perspective for future work.
Acknowledgements:
J.-P.M was partially supported by ANID (Chile) via the FONDECYT grant #1201894 during the first stages of this work. He also thanks the France 2030 framework programme Centre Henri Lebesgue ANR-11-LABX-0020-01 for creating an attractive mathematical environment.
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work the author used ChatGPT in order to smoothen the expression in English. After using this tool/service, the author reviewed and edited the content as needed and takes full responsibility for the content of the publication.
Data availability statement
No dataset is related to this work.